### v3

# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

Abstract.
Decentralized Finance (DeFi) incidents stemming from the exploitation of smart contract vulnerabilities have culminated in financial damages exceeding $3$ billion US dollars. Existing defense mechanisms typically focus on detecting and reacting to malicious transactions executed by attackers that target victim contracts. However, with the emergence of private transaction pools where transactions are sent directly to miners without first appearing in public mempools, current detection tools face significant challenges in identifying attack activities effectively.
Based on the fact that most attack logic rely on deploying one or more intermediate smart contracts as supporting components to the exploitation of victim contracts, in this paper, we propose a new direction for detecting DeFi attacks that focuses on identifying adversarial contracts instead of adversarial transactions. Our approach allows us to leverage common attack patterns, code semantics and intrinsic characteristics found in malicious smart contracts to build the LookAhead system based on Machine Learning (ML) classifiers and a transformer model that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions of potential zero-day attacks. Our contributions are three-fold: First, we construct a comprehensive dataset consisting of features extracted and constructed from $210,643$ recent contracts and $375$ hand-picked then carefully reviewed and identified adversarial contracts deployed on the Ethereum and Binance Smart Chain (BSC) blockchains. Secondly, we design a condensed representation of smart contract programs called Pruned Semantic-Control Flow Tokenization (PSCFT) and use it to train a combination of ML models that understand the behaviour of malicious codes based on function calls, control flows and other pattern-conforming features. Lastly, we provide the complete implementation of LookAhead and the evaluation of its performance metrics for detecting adversarial contracts. Our experiments show that LookAhead achieves an F1-score as high as $0.8966$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta, with a False Positive Rate (FPR) at only $0.16\%$.
copyright: none
1. Introduction
Decentralized Finance (DeFi) has gained significant traction within the blockchain ecosystem over the recent years and is continuing to evolve rapidly, stimulating the rise of a diverse array of DeFi applications such as: decentralized exchanges (DEXs) (Adams et al., [n. d.]; Martinelli and Mushegian, [n. d.]), lending platforms (AAV, 2024; Robert and Geoffrey, [n. d.]), and derivatives instruments (Juliano, [n. d.]). The total value locked (TVL) in these DeFi protocols has exceeded \$100 billion in mid 2024 (Def, 2024). Among the various blockchain networks, Ethereum (Wood et al., 2014) and Ethereum Virtual Machine (EVM)-compatible chains (Arb, 2024; BNB, 2024; Opt, 2024; Pol, 2024) have formed a particularly thriving DeFi ecosystem, collectively they play a pivotal role in shaping the future of the blockchain landscape. At the heart of Decentralized Finance are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. The functionalities of smart contracts are invoked by users through transactions, after which the blockchain consensus algorithms ensure predefined actions and state transitions are execute correctly and autonomously on-chain. However, just like traditional software, smart contracts are fundamentally code-based programs, which are susceptible to vulnerabilities and various other security risks. In fact, DeFi incidents resulting from smart contract attacks have caused over 3 billion US dollars in financial losses (DeF, 2024; Lea, 2024; Zhou et al., 2023).
DeFi Attacks. Targeting smart contracts, there are mainly two types of attack patterns commonly used by DeFi attackers: an attacker can either directly invoke specific public functions of the victim contracts in order to exploit a vulnerability (Tre, 2024; Sus, 2024b), or they can deploy an adversarial contract containing the complete attack logic onto the blockchain first, then initiate the attack through the intermediate contract by calling their entry functions afterwards (Bea, 2024; Bur, 2024; Pan, 2024; Sus, 2024a).
Limitations of Existing Work. Existing real-time defense mechanisms leverage heuristics (Qin et al., 2023b; Wu et al., 2023; Zhang et al., 2020; Zhou et al., 2020) and Machine Learning (ML) (Gai et al., 2023; Wang et al., 2022) techniques to detect and respond to yet-to-be-confirmed transactions in public mempools that are adversarial in nature. The main issue with this transaction-based detection approach is that it faces significant challenges in identifying private adversarial transactions. Namely, attackers can utilize private mempool services to send adversarial transactions directly to miners, evading visibility from other participants of the blockchain network before the transaction is confirmed and effectively bypassing attack detection. The ability to send private transactions was originally developed by researchers and industry practitioners to counter malicious activities on-chain such as front-running, but such services have unfortunately been increasingly abused by attackers for vulnerability exploitation purposes.
Figure 1. The proportion of private transactions
in DeFi attacks between 2022 and 2023.
Figure 2. Cumulative distribution function (CDF) of the rescue time frame in DeFi attacks.
Key Observations. Based on our empirical study of the historical DeFi incidents over the past few years, we found that out of the $142$ DeFi attacks on Ethereum collected by us, $80$ cases ($56.3\%$) involve transactions using private mempool services. Moreover, as shown in Figure 2, the proportion of adversarial transactions sent using private services exhibits a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023. We emphasize the critical importance of establishing a new approach for detecting DeFi attacks that can remain effective even when private mempool services are employed by attackers. Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor function during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims. As illustrated in Figure 2, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$. Based on the aforementioned studies and observations, we design our method for preventing DeFi attacks via unveiling adversarial contracts.
Our method. In this paper, we propose the LookAhead system, representing a new direction for detecting DeFi attacks, i.e., focusing on identifying adversarial contracts instead of adversarial transactions, and allowing us to proactively identify potential attack intentions and victims, even in the case of private adversarial transactions.
By shifting the focus to analyzing and identifying adversarial contracts, we filled several gaps currently present in traditional transaction-based detection tools. We first make an empirical observation that adversarial contracts used in DeFi attacks often contain similar patterns, for example, over 70% of the contracts in our adversarial contract dataset use funds from anonymous sources for deployment; more than 98% are closed-sourced, etc. We also note that depending on the vulnerabilities targeted, attackers write contract code in specific manners. For instance, flashloan and reentrancy attacks are two of the most notorious and frequently studied exploitation. Both of them require the establishment of complex function call chains and the implementation of attack logic within callback functions. To use these behaviours to our advantage for identifying adversarial contracts, we using the stacking technique to design three ML models, one transformer, multiple candidate classifiers, and a meta classifier that, when used together, can understand the intrinsic characteristics of malicious smart contracts based on code semantics and common patterns. To support the ML models in understanding the behaviour of Solidity smart contract code, we develop a specialized tokenization method called Pruned Semantic-Control Flow Tokenization (PSCFT) that is integrated directly into a custom contract bytecode decompiler. By leveraging modern ML architecture and models, we can produce batch predictions for newly created contracts on-chain in a timely manner with amortized $t^{pred}\ll 100s$. The result of our design is LookAhead, a system that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions of potential zero-day attacks.
To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our attack contract dataset. And in total $796,437$ contracts deployed on Ethereum from June 2022 to June 2024 were collected, with $210,643$ of them being used as benign samples after filtration. We extract a selection of features from the contracts capturing patterns exhibited during both their implementation and their deployment stages. Given that our dataset contains a combination of contracts from both recent and historical attacks, we believe that our feature selection will remain reasonably relevant and effective even in the rapid growing landscape of decentralized finance. The selection methods and criteria for building this dataset are described in details in section 4. These data have been made publicly available to facilitate future research in this domain.
We build our classifiers using multiple supervised ML algorithms and conduct a comprehensive performance evaluation. Experiment results show that our method for identifying adversarial contracts performs exceptionally well with an F1-Score produced by the KNN meta classifier reaching as high as $0.8966$, and a false positive rate of only $0.16\%$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta.
Contributions. We summarize our contributions as follows:
  • We propose a new direction for detecting DeFi attacks that focuses on identifying adversarial contracts instead of adversarial transactions.
  • A comprehensive dataset of adversarial and benign smart contracts with a variety of features pre-extracted, pre-processed and attached to each item.
  • A custom designed smart contract bytecode lifting and analysis pipeline with integrated semantic tokenization protocol that is able to efficiently extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
  • An ML-based system consisting of two open-weights classifiers and one meta classifier model that is pre-trained, fine-tunable, and capable of conducting accurate and just-in-time detection of potential adversarial attacks, even when private adversarial transactions are employed by attackers.
  • The implementation and a detailed evaluation as well as performance analysis of LookAhead’s detection system and methods.
After providing some background in Section 2, we present an overview of LookAhead in Section 3. Section 4 details our dataset building strategies and Section 5 explains our methods, Section 6 reports our experimental results. After a brief discussion in Section 7 and 8, we conclude the paper in Section 9.
2. Background
2.1. Ethereum & Decentralized Finance (DeFi)
Blockchain is a distributed ledger built on top of a peer-to-peer (P2P) network, maintained through consensus mechanisms to ensure the consistency of the ledger. Ethereum (Wood et al., 2014) is one of the most used blockchain platforms. It was the first to introduce the Ethereum Virtual Machine (EVM) that support smart contracts, and has inspired the development of a series of EVM-compatible chains. DeFi is a blockchain-based financial ecosystem powered by smart contracts to offer financial services in a more open and transparent manner. There are several unique concepts in DeFi that differ from traditional finance.
Accounts. Accounts are entities that can hold the Ether token and initiate transactions. Ethereum has two types of accounts: External Owned Accounts (EOAs) and Contract Accounts (CAs). EOAs are controlled by users holding private keys, while CAs are containers of contract code and storage.
Transactions. Transactions are used to transfer Ether, invoke contracts, or create contracts. After a transaction is sent, it is first placed in the mempool of a blockchain node. Public mempools are transparent and openly accessible and filled with pending transactions, waiting to be selected and finalized by block producers.
Smart Contracts. Smart contracts are instantiated objects stored on the blockchain. They are typically written in languages like Solidity (Dannen, 2017) and compiled into bytecode that can be executed within the EVM, a stack-based virtual machine supporting Turing-complete instructions.
Flashloan. Flashloan is a unique form of lending that leverages the atomic nature of blockchain transactions to allow users to borrow and repay cryptocurrency within a single transaction. This mechanism lets users temporarily possess large amounts of tokens for a small fee. Despite of its convenience, it has been exploited in numerous DeFi attacks (Cao et al., 2021; Qin et al., 2021; Wang et al., 2020).
2.2. Maximal Extractable Value (MEV)
Traditionally, miner determine the order of transactions in a block based on gas prices. Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the gas price, allowing for front-running (Eskandari et al., 2020; Torres et al., 2021) or back-running to happen. The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Maximal Extractable Value (MEV) (Daian et al., 2020). It has been proven to have a significant impact on the security of blockchain networks (Qin et al., 2022).
Front-running. Front-running is a fundamental means used to extract MEV, typically employed in arbitrage, liquidation, and sandwich attacks. Front-runners monitor the public mempool for victim transactions that meets specific conditions, and attempt to raise the gas price to execute their own transactions first against the original transaction. Furthermore, front-running has also been used in intrusion prevention systems (Qin et al., 2023a; Zhang et al., 2023), as a way to front-run adversarial transactions in order to prevent DeFi attacks.
Private Mempool Services. In response to MEV, researchers and industry practitioners have proposed various countermeasures (Yang et al., 2022). The MEV-auction solution introduced by Flashbots (Fla, 2024) has recently gained popularity (Weintraub et al., 2022). This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others in the public mempool. However, such private mempool service are also utilized by attackers to send private adversarial transactions, aiming to execute attack logic and effectively bypass various detection methods.
3. Overview of LookAhead
3.1. Motivating Example
Figure 3. Read-only reentrancy attack on Conic Finance, using a private adversarial transaction to evade detection, could have been prevented during the rescue time frame.
On July 21, 2023, DeFi protocol Conic Finance suffered a major exploit (DLNews, 2024) causing financial losses exceeding \$3 million. We present an overview of the attack process in Figure 3. Notably, the adversary transaction that triggered the attack was labeled as an MEV Transaction on Etherscan  [^1] https://etherscan.io/tx/0x8b74995d1d61d3d7547575649136b8765acb22882960f0636941c44ec7bbe146, indicating the use of private mempool services. This allowed the attacker to bypass the public mempool and invalidate any potential front-running defense attempts. This attack highlights the limitations of traditional attack prevention mechanisms that rely on detecting adversarial transactions, particularly when private transactions are used.
3.1.1. Rescue Possibility
Despite the attacker using private mempool services to conceal the adversarial transaction, there was still an opportunity to prevent the attack, based on the following facts: 1) The ConicEthPool contract’s address — the vulnerable target — was written in plaintext in the adversary contract at the time of contract deployment; 2) There was a 36-second window (3 blocks) between the deployment of the adversarial contract and the execution of the adversarial transaction; 3) The ConicEthPool contract had an access-controlled shutdownPool function, which offers the ability to perform an emergency pause to block the adversary’s attempt to call the deposit function. Thus, detecting the adversarial contract at the deployment stage, before the confirmation of the adversarial transaction, could have prevented the exploit, regardless of whether the attacker used a private mempool service or not.
3.1.2. Adversarial Patterns
While DeFi attack logic can vary based on in the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives. Our analysis of numerous adversarial contracts reveals a clustered set of patterns including the use of anonymous fund sources, closed-source code, and frequent token-related function calls. Additionally, adversarial contracts often have distinctive interaction logic with victim contracts, which differs significantly from the patterns observed in normal contracts, resulting in easily distinguishable code structures and external call flows. Based on these observations, ML techniques can be employed to effectively identify adversarial contracts — regardless of the specific attack type — by learning and predicting these common patterns.
3.2. Threat Model
3.2.1. DeFi Attacks
There are many types of security breaches with complicated root causes that could lure attackers into exploiting a victim contract. In different attack scenarios, an adversary could use different techniques to conduct their exploitation. In this section, we present these attack methods, define key concepts involved in the attack process, and restrict LookAhead’s detection scope to a certain area of focus.
Figure 4. a) Left: attack directly, b) Right: attack through adversarial contracts.
Attack Methods. The attack process of two common attack methods are depicted in Figure 4.
  1. 1)
    Attacking directly via adversarial transactions. In the presence of complex attacking logic involving multiple transactions, the attackers often bear the risk of encountering potential half-way failure. A half-way failure arises when only part of an attack was executed successfully and the rest fail to proceed due to various reasons (e.g., transaction conflicts).
  2. 2)
    Attacking through adversarial contracts. This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failures thanks to the atomic nature of a blockchain transaction. Using this method, an attacker first deploys a smart contract on-chain containing the core attack logic, then initiate the attack by calling the entry point function of the adversarial contract. Most DeFi attacks follow this pattern, hence we focus on addressing this type of attacks.
Scope of Detection. We specify three attack scenarios that fall outside of the scope this study. 1) Attacks that do not require contract deployments. Those resulting from private key leakage or attackers directly invoking public functions of the victim contracts using EOAs are outside the scope of our detection. 2) A small portion of attacks are completed directly through the constructor function at the time of contract deployment. Since the adversarial contracts used in these attacks already exhibit execution traces of the actual attack at the time of deployment, we consider them as adversarial transactions, which fall outside the scope of our detection. 3) We do not place our attention on detecting attacks such as rug pulls as they can only be launched by privileged users.
LookAhead focuses on unveiling attacks with intention to cause financial loss for DeFi protocols by exploiting vulnerabilities in on-chain smart contracts. To identify an attack, our goal is to detect adversarial contracts that contain the core attack logic. Other contracts, such as those used solely for concealing profits to avoid detection, fake token contracts to facilitate the main attack, and counterattack contracts used by front-running bots will not be included in our analysis.
3.2.2. Adversaries
We consider a resource bounded adversary $\mathcal{A}$ that is financially rational, capable of deploying an arbitrary amount of smart contracts and execute both public and private transactions on EVM-compatible blockchains.
Attacker Assumptions. Based on previous research findings (Zhou et al., 2023), our study makes the following assumption about $\mathcal{A}$:
  1. 1)
    $\mathcal{A}$ do not initiate attacks by batching $tx_{deploy}$ and $tx_{first}$ together in an atomic transaction during contract deployment, meaning that $\mathcal{A}$ splits an attack into two stages following the second pattern described in Figure 4;
In a prior research (Parhizkari et al., 2023), it was observed that 79.49% of attacks in their dataset exposed the victim contract address before the adversarial transaction. Using the same method, we found 202 attacks (53.9%) with similar exposure behaviour in our dataset of $375$ adversarial contracts. Hence, we make one additional assumption about $\mathcal{A}$:
  1. 2)
    $\mathcal{A}$ hard-code exploitation targets’ addresses in their malicious contracts whenever possible, and the victim’s address will emerge before the first attack transaction $tx_{first}$.
3.2.3. Victim of Attacks
We consider those contracts which contain vulnerabilities that have been exploited by attackers and suffered a financial loss the victims of a DeFi attack.
3.2.4. Rescue Time Frame
Our threat model leaves a short rescue time frame for detection and prevention to take place by the defenders and the targets, and enables the possibility of on-time reaction by victims of ongoing attacks. The rescue window is defined by $t^{window}=t_{first}-t_{deploy}-t^{pred}$, it has a typical value observed in our dataset of $t^{window}\geq 100$. Within the rescue time frame several counterattack actions could be taken, we describe those in more details in §3.3.
3.3. System Model
In order for the LookAhead system to provide just-in-time identification of adversarial contracts as they are deployed onto the blockchain, a streamlined process consisting of a series of components is used. We show an overview of our system in Figure 5.
Chain Monitoring. Through a blockchain network node or third-party node monitoring services, we monitor on-chain data in real-time, and acquire the deployed contracts and their basic information such as deployer, input data, gas used, bytecode and so on whenever needed. Via chain explorers such as Etherscan, we are able to retrieve the verification status of contracts as well. These data will be combined and fed into a classifier as features for ML inference at a later stage.
Figure 5. An overview of the LookAhead system.
Binary Lifting. We do not use any source code information from the contracts, but only the bytecode deployed on the blockchain to ensure our model is capable of analyzing any smart contract on EVM blockchains. To do that, LookAhead uses the open-source binary lifting and analysis library Gigahorse (Grech et al., 2019) to decompile the contract bytecode and generate intermeidate representation (IR). We then run a program to visualize the Control Flow Graph (CFG) based on the IR.
PSCFT and Feature Extraction. With the basic CFG, we use a specialized feature extraction program to pull out features including but not limited to: token-related call information, flashloan usage information, etc. Finally, we convert the CFG into a condensed text form (PSCFT) and populate it with semantic information such as function names and external call target address labels.
Classifier Inference. Input features are properly tokenized and formatted before they run through our trained classifiers. These ML models are made up of a transformer, an optimal candidate classifier, and a meta classifier trained using supervised machine learning methods for binary classification. They are able to output either the classification result $label_{pred}\in\{0,1\}$ indicating whether the contract is adversary ($1$) or not (00), or they can produce a confidence score $0\%\leq p_{pred}\leq 100\%$ representing the possibility of the input contract being malicious.
Defense Mechanisms. After an ongoing attack has been identified by LookAhead, there are several defense techniques that can potentially be used by defenders and victim contract owners. 1) If the victims’ addresses were found in the adversarial contract, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts. Moreover, if contract owners are willing to give us or any LookAhead system operator admin permissions to execute their shutdown methods on behalf of them, we can automatically react to possible attacks in an even more timely manner. 2) We plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses. If enough protocol users and developers incorporate our service as part of their access control pipeline, they can take full advantage of the real-time protection abilities of LookAhead at minimal to no cost to their operations.
4. Dataset Building
4.1. Contract Collection
To perform supervised training of our classifier models, we collect a large set of past and recent smart contracts. For the positive (adversarial) dataset, we curated a collection of adversarial contracts used in past real-world DeFi incidents. For the negative (benign) dataset, we follow a methodical process to choose samples that are very unlikely to be related to any attacks.
4.1.1. Adversarial Contracts
To collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as GitHub repositories documenting DeFi incidents (DeF, 2024; Lea, 2024), and alerts from security companies (blo, 2024; cer, 2024; SlowMist, [n. d.]). We conduct a meticulous review of these incidents to identify adversarial contracts used by attackers. Between April 2020 and July 2024, we collected 506 DeFi attacks. Of these, $29$ attacks ($5.7\%$) involved direct attacks from EOAs without using contracts, and $37$ attacks ($7.3\%$) were executed directly within the contract constructor. Both cases fall outside the scope of our detection. While attacks spanned multiple blockchains, the majority were on Ethereum and BSC. To simplify the study, we exclude 65 incidents happened on other chains. This left us with $375$ adversarial contracts: $142$ from Ethereum and $233$ from BSC. Following prior research (Zhou et al., 2023), we classify the attacks into three categories: Untrusted or unsafe calls ($13.3\%$), Access control mistakes ($14.4\%$), and Coding mistake ($72.3\%$). The first category includes issues like reentrancy, the second covers access control flaws, and the third involves coding mistakes such as arithmetic errors and logic absences.
4.1.2. Benign Contracts
To construct benign samples, we first collected a total of $796,437$ contracts deployed between June 2022 and June 2024 using Google’s BigQuery service. And we analyzed the transaction history of these contracts, focusing on the number of unique addresses interacted with the contracts from June 2022 to August 2024. The result showed that the adversarial contracts in our dataset often have very few unique users, with $355$ of them ($94.7\%$) interacting with only a single address, namely the attacker’s account, and a maximum of $6$. This low interaction count is consistent with the behavior of adversarial contracts, which typically feature access control and are designed for single-use only to attack specific victims. Based on this observation, we select contracts with at least $10$ unique interacting addresses as the determining factor for their benign nature. This filtering resulted in $210,643$ benign contracts.
4.2. Feature Selection
Based on empirical observations, we construct a selection of features for use in LookAhead’s ML model training. The set consists of features spanning over both the implementation and the deployment stages of a DeFi attack, aiming to capture any potential unexpected patterns of behaviour that belong to attackers. In this section, we explain our rationale behind the set construction.
4.2.1. Deployment Features
A typical workflow for deploying a contract to a blockchain consists of 1) Sourcing the fund to cover deployment costs, 2) Sending a transaction containing the contract bytecode and other related data on-chain, 3) Verifying the source code of the contract to demonstrate its security and trustworthiness to the wider community. During the process, a malicious individual often behave very differently from a normal one, we use this to our advantage by selecting several measurable properties as supporting factors to help us make such a distinction.
Creator Attributes: $\{nonce,fund\_source\}$. $nonce$ indicates number of transactions sent from the contract creator before the deployment of the contract. A higher $nonce$ suggest a more active EOA address. We are also interested in the $source$ from which the contract creator initially obtained their funds. To conceal their identities, attackers typically fund their attacks from a mixer (e.g, Tornado Cash (Wang et al., 2023)) or an instant exchange (Cha, 2024; Sid, 2024) without sign-up or KYC (Know Your Customer) requirements, then transfer the funds to a new account for subsequent transactions. In contrast, regular users are more likely to source their funds from KYC-required centralized exchanges.
We categorize fund sources into four types based on address information:
  1. 1)
    Safe: Fund sources with relatively secure origins, such as KYC-required centralized exchanges, which are less likely to be used by attackers.
  2. 2)
    Anonymous: Fund sources with anonymous origins, including mixers and some KYC-free exchanges, which are more susceptible to be used for malicious purposes.
  3. 3)
    Bridge: Funds originating from cross-chain transfers. The source could be anonymous, hence they are considered potentially usable for attacks.
  4. 4)
    Unknown: Fallback label for those that fail to fall into the above categories.
Based on this taxonomy, we conduct a statistical analysis on the fund source of our contract samples. For our benign samples, more than three-quarter of them have Safe sources, approximately one-tenth of their labels are Anonymous. In contrast, the majority of the adversarial contracts have Anonymous fund sources, with only a small proportion having Safe labels.
Transaction Data: $\{value,input\_data\_length,gas\_used\}$. Depending on the contract’s design, the contract creator has the ability to transfer native tokens to the contract upon deployment. The amount transferred is indicated by the $value$ attribute associated with the transaction. Adversaries typically do not perform such a transfer. The input data field of a deployment transaction holds the contract’s bytecode, we measure its length to account for the size of the contract. Additionally, since the gas consumption of the deployment transaction is proportional to the complexity of the contract, we combine these metrics to effectively measure the scale of the contract.
Verification Status: $\{verified\}$. The $verified$ status can be granted to contracts if their source code and compiler settings are submitted to platforms like Etherscan. Verified contracts are open-source, anyone can view their code. Many contract development frameworks, such as Foundry (fou, 2024), offer options to verify contracts upon deployment. It has become increasingly common for developers to verify their contracts to demonstrate security and trustworthiness. Attackers, on the other hand, typically do not have the incentives to do so since it would expose their malicious intentions.
4.2.2. Implementation Features
Based on empirical observations, some of the characteristics that can be used to determine whether a contract is malicious or not include: implemented functions, internal and external function calls. To capture other hidden or intrinsic properties of a contract, we also design a special technique called Pruned Semantic-Control Flow Tokenization (PSCFT) that provides a condensed representation of the behaviour of smart contracts.
Implemented Functions: $\{func\_count,flashloan\_callback\_count\}$. Since adversarial contracts typically do not involve complex business logic, they often implement fewer functions. In our dataset, benign contracts frequently incorporate $\geq 20$ public functions, while adversarial ones tend to implement much fewer. We also calculate the number of flashloan callback functions and the proportion they take up over all public functions. This choice is motivated by the fact that many attack scenarios require attackers to have significant token holdings, which most adversaries do not naturally possess. Consequently, they resort to utilizing flashloan to temporarily borrow a substantial amount of tokens and execute attack logic within corresponding callback functions. Since the size of lending pools are limited, attackers often borrow tokens from multiple sources, resulting in multiple flashloan callbacks. In our dataset, more than 60% of adversarial contracts implement flashloan callbacks, while almost all benign contracts do not involve such callbacks.
Function Calls: $\{token\_call\_count,max\_token\_call\_count,avg\_token\_call\_count,delegate\_c\\ all\_count,selfdestruct\_count\}$. We take into consideration external calls related to ERC-20 tokens, including their count and proportion. Generally, the objective of DeFi attacks is to gain financial benefits, i.e., acquire valuable tokens. Therefore, external calls within adversarial contracts often involve token-related functions. The token-related functions we selected include, but are not limited to, common token standard (ERC, 2024a, b) functions and popular DEX (Uniswap, 2024; Adams et al., [n. d.]) functions. In our dataset, over $90\%$ of benign contracts do not have token-related calls. In contrast, almost all adversarial contracts involve token-related calls, with over two-thirds of them even having $10$ or more.
Furthermore, we introduce two additional features to assess the behaviour of public functions in a smart contract. $max\_token\_call\_count$ is designed to capture the maximum number of token-related interactions within the execution flow of any public function of a contract, which helps us identify contracts with exceptionally high token interactions. Moreover, $avg\_token\_call\_count$ is also measured to demonstrate the average number of token-related calls in contracts.
We also include the number of delegatecall instructions, as these are often implemented in proxy contracts. Additionally, some attackers have a tendency to self-destruct their contracts after completing the attack, so we also consider the presence of the selfdestruct instruction.
Code Semantics: $\{PSCFT\}$. To understand the behaviours exhibited by DeFi attack adversaries and capture features that are potentially hidden from our empirical observations, we design a new representation for smart contracts called Pruned Semantic-Control Flow Tokenization ($PSCFT$). PSCFT improves upon the intermediate representation (IR) generated by Gigahorse to produce richer contextual information about the implementation of a contract. After constructing CFGs based on IR and going through an augmentation pipeline that retrieves and replaces function signatures and external addresses with appropriate semantic labels in the CFGs, PSCFT can effectively present a textual summary of function canonical names, control-flow graphs and external call chains. Figure 6.3 shows a contract PSCFT excerpt. One of the benefits of using PSCFT over IR is that IR alone does not provide useful information such as call target labels, and the size of the tokenization vocabulary is less manageable, rendering it not ideal for NLP training.
5. Methodology
Figure 6. LookAhead’s system building workflow
In this section, we present our methodology for building the LookAhead system that performs adversarial contract classification. An overview of the complete framework is depicted in Figure 6.
5.1. External Data Retrieval
Some of the features used in our dataset are obtained from external sources, we identify related third-party services in the following section. Based the data, we develop algorithms such as fund source tracing to retrieve specific fields to be used for both training and inference of our ML models.
Transaction Data. In this study, we focus on investigating and demonstrating the viability and effectiveness of the LookAhead system. Therefore, for the sake of simplicity, instead of running our own blockchain nodes, we obtain transaction raw data from an RPC (Remote Procedure Call) provider such as Alchemy (Alc, 2024) that relays our requests directly to other nodes in the blockchain network. Prior to the practical deployment of LookAhead, we recommend setting up dedicated nodes for minimized data processing delays. We store and format contract deployment transactions and their associated attributes such as nonce, value, and gas consumption, along with the sender and input data properties for use in subsequent phases of our feature extraction pipeline.
Chain Explorer Data. We leverage the APIs provided by chain explorers, Etherscan and BscScan to retrieve information about the verification status of contracts. However, we cannot obtain the exact time for contract verification. The effectiveness of this feature is based on the assumption that developers verify their contracts upon deployment (with minimal to no delay). We describe the potential impact that this assumption could bring to the performance of LookAhead in §6.2.
Furthermore, we develop a fund source labeling algorithm that recursively follow the transaction flow and trace all the way back to the original fund holder. Utilizing the address label cloud (lab, 2024) and public name tag provided by chain explorers, we manually review and assign fund source labels to each group of addresses. As an example, Tornado.Cash was given the label of Anonymous, while centralized exchanges like Coinbase were given the label of Safe.
5.2. Pruned Semantic-Control Flow Tokenization (PSCFT)
Based on state-of-the-art EVM contract binary lifter and analysis framework Gigahorse (Grech et al., 2019), we design a pipeline for 1) Constructing intra-contract Control Flow Graphs (CFG) and external call chains based on Intermediate Representation (IR), 2) Aggregating and extracting dataset features from EVM bytecode, and 3) Generating textual summaries for the smart contracts in our dataset.
5.2.1. Implementation Feature Extraction
To enable the extraction of implementation features (cf. §4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse. First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and $selfdestruct\_count$. Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers) in combination with the PublicFunctionSelector relation, which maps public functions to their selectors, to determine the total number of public and private functions. This combination also helps identify flashloan callback functions. Finally, to analyze the external function calls within a contract, we design a rule called ExternalCallResolved. This rule provides details about all resolved external calls made within the contract, including their selectors and signatures. Additionally, we implemented logic to handle special cases, such as identifying fallback functions and transfers of native tokens, based on the arguments and values passed to the calls. With the help of ExternalCallResolved, we can keep track of token-related calls and their related information including their total count, maximum occurrences, and average usage.
5.2.2. CFG Construction and Pruning
Based on the IR of a contract, we construct a function-by-function visualization $F=\{F^{private}\cup F^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements. Functions have 4-byte signatures derived from the Keccak hash (SHA3) of their canonical names and can be either private or public. Each function contains multiple basic blocks (BB) chained together ($\{\mathcal{F}_{i}=\{BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|\mathcal{F}_{i}|}\}\}$), and each block contains multiple statements. Blocks and statements are identifiable via unique numerical IDs, for example, $90x11b$ 9. The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B=\{\mathcal{BB}^{i}_{j}:\mathcal{BB}^{i}_{j}\in\mathcal{F}_{i}\}$ where $\mathcal{F}_{i}\in F$, and $(\mathcal{BB}^{i}_{x},\mathcal{BB}^{i}_{y})$ is an arc in $A_{i}$ if $\mathcal{BB}^{i}_{x}$ is a predecessor of $\mathcal{BB}^{i}_{y}$ and $\mathcal{BB}^{i}_{y}$ is a successor of $\mathcal{BB}^{i}_{x}$. Statements could be constant/variable declarations, internal/external function calls, function terminations, memory copies, storage initialization, etc., we focus on constructing a CFG based on function call statements.
The pruning process starts with the sorting and reordering of functions based on function types (public first, then private) and their canonical names. Gigahorse can only recover the names for some of the public functions and external call target methods, hence if the derivation of canonical names from function signatures failed for some public function $\mathcal{F}^{public}_{i}\in F^{public}$, then $\mathcal{F}^{public}_{i}$ will be given a name of $UnknownFunction_{i}$. For any private function $\mathcal{F}^{private}_{j}\in F^{private}$, we give it a canonical name of $InternalFunction_{i}$. Next, we remove all statements not related to function calls from every block $\mathcal{BB}^{i}_{j}\in B_{i}$. Then, we perform a depth-first search (DFS) over $G$ to remove any block $\mathcal{BB}^{i}_{k}$ that contains no statements (i.e. $|\mathcal{BB}^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours. Finally, we assign canonical names $BB^{i}_{j},j\in[0,|B|-1]$ to basic blocks that are still left in $B_{i}$.
5.2.3. Semantic Recovery
For an external call statement like the one shown in section PSCFT and Feature Construction in Figure 6, we aim to recover: 1) An address label for 0xBA12222222228d8Ba445958a
75a0704d566BF2C8
, 2) A human-readable name for the target method 0x5c38449e.
Similar to our fund source labelling algorithm, we attempt to match the target address with an address associated with a known label in Etherscan’s address label cloud and public name tags. For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label. Contract developers might choose to store the target address in the EVM storage or as a variable/constant which affects our ability to directly perform address matching. We use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage. If this process did not result in a matching label, we assign the address a default label UnknownTarget.
Function signatures are generated by hashing the function names with a secure hashing algorithm, and the hashing process is a one-way trap door. However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings. Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their Application Programming Interfaces (API) upon failed recovery by Gigahorse.
If the aforementioned process was completed successfully, the external call statement will be converted to Balancer.flashLoan(…args). And we can produce a textual summary for smart contracts containing control-flow and semantic call information simply by concatenating $\mathcal{F}_{i}=\{\mathcal{BB}^{i}_{0},\cdots,\mathcal{BB}^{i}_{|\mathcal{F}_{i}-1|}\}\in F$ together in the form shown in Figure 6.3, Semantic Recovery.
5.3. Classifier Design
5.3.1. Semantic Transformer
We train a transformer-based Natural Language Processing (NLP) model $\mathcal{M}^{transformer}$ solely on the PSCFT constructed from the contracts in our dataset to learn the semantics and behaviours of the implementation of both benign and adversarial smart contracts. Following the standard architecture of transformers (Vaswani, 2017), we design:
  1. 1)
    A word2vec (Mikolov, 2013) embedding layer for positional encoding of PSCFT syntax and semantics,
  2. 2)
    Multiple encoder layers chained together, each made up of a multi-head attention layer followed by a feed-forward layer,
  3. 3)
    Pooling and dense layers for outputting classification results.
To ensure the model’s generalization ability and improve training efficiency, we implement regularization techniques such as early-stopping and dropouts to prevent model over-fitting. Overall, this transformer model provides a semantic view of DeFi smart contracts.
5.3.2. Ensemble Classifier
Aiming to embed the features selected based on our empirical observations into the semantic view of the transformer model, we design an ensemble classifier that merges the classifier models trained on basic features with the transformer model trained on PSCFT together to produce a meta classifier model with better predictive performance.
Candidate Classifiers. Once all features have been extracted, we can proceed to train a supervised learning classifier using a variety of ML algorithms, outputting multiple candidate models for further evaluation. These models include classical approaches such as Logistic Regression $\mathcal{M}^{classifier}_{LR}$ and Decision Trees $\mathcal{M}^{classifier}_{DT}$, ensemble learning methods such as bagging-based Random Forest $\mathcal{M}^{classifier}_{RF}$ and XGBoost $\mathcal{M}^{classifier}_{XGBM}$ based on boosting.
Due to the limited quantity of adversarial samples, to mitigate the bias introduced by class imbalance, we employ the ADASYN oversampling method (He et al., 2008) that synthesizes additional samples to bring the number of adversarial samples closer to that of the benign ones in the training dataset.
For the fund source feature, which consists of categorical data (ref. §4.2.1) with low cardinality, we apply one-hot encoding. To facilitate better comparability among input values in numerical terms of varying dimensions, we standardize the dataset using Z-score normalization. This process brings each feature value around to the centre of 00 with a standard deviation of $1$.
After training has completed for each of the candidate classifiers, we evaluation their performance using metrics such as Recall and F1-Score to determine an optimal model $\mathcal{M}^{classifier}_{optimal}\in\{\mathcal{M}^{classifier}_{LR},\cdots,\mathcal{M}^{classifier}_{XGBM}\}$ for the meta classifier training and inference in the next stage.
Meta Classifier. Using the stacked generalization (stacking) machine learning algorithm, we design a blender that stacks $\mathcal{M}^{transformer}$ and $\mathcal{M}^{classifier}_{optimal}$ together to form a higher level model $\mathcal{M}$ known as a meta classifier. By feeding the predictions produced by the two base models into the meta model, we are able to 1) Combine the strengths of each model and automatically learn the optimal way to construct the inference result based on model significance, 2) Optimize the final classification boundaries to achieve more accurate detection of adversarial contracts.
We use K-Nearest Neighbors (KNN) as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space. KNN is well-suited for combining the outputs of different classifiers such as the transformer model and one of the candidate classifiers, because it is able to combine the complementary strengths of both models by analyzing their predictions and finding the optimal decision boundary.
6. Evaluation
We implement LookAhead in Python using open source machine learning frameworks PyTorch and Tensorflow to achieve both CPU and GPU-accelerated training and inference of our classifier models. All experiments are conducted on a server running a 3.70GHz Intel Xeon processor with 32 threads and a NVIDIA A100 GPU with 40GB of VRAM.
6.1. Experiment Setup
6.1.1. Dataset
Token and proxy contracts are two of the primary types of contracts used on blockchains, while they are commonly used by developers, most adversaries do not rely on those types of contracts for their core attack logic. To identify contract types, we perform a simple check based on PSCFT to determine whether a contract follows known standards such as ERC20 and ERC1967. By analyzing whether the contracts match common standards, we found that in the dataset we built, $184,386$ benign samples were token contracts, $10,037$ were proxy contracts, and none of the adversary contracts belonged to these two categories. Therefore, we remove them from both our training and testing dataset the same way as it was in previous system Forta (For, 2024).
A distinguishing aspect of our research when compared to traditional supervised machine learning problems is the explicit chronological order of contract deployment. Contracts deployed over a different periods of time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve and change over time as well. One of LookAhead’s objectives is to provide a ML classifier model that is pre-trained on past DeFi incidents but capable of detecting potential future attacks. Consequently, traditional cross-validation techniques, which assume input values are independent with respect to each other, are not suitable for our evaluation. Instead, to achieve such an objective, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing. Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
6.1.2. Research Questions
Model Accuracy Precision Recall F1-Score FPR
Candidate Classifiers LR 0.9925 0.8788 0.7838 0.8286 0.0025
DT 0.9838 0.6222 0.7568 0.6829 0.0108
RF 0.9913 0.8594 0.7432 0.7971 0.0029
XGBoost 0.9932 0.8333 0.8784 0.8553 0.0041
NLP Classifier Transformer 0.9876 0.7429 0.7027 0.7222 0.0057
Forta - 0.8778 0.5520 0.6208 -
Meta Classifier KNN 0.9953 0.9286 0.8667 0.8966 0.0016
Table 1. Performance breakdown of different models.
We attempt to address the following research questions:
  1. 1)
    RQ1. How effective is LookAhead’s system of classifiers in detecting the adversarial contracts used in DeFi attacks, especially when compared to previous state-of-the-art solutions?
  2. 2)
    RQ2. How efficient can the LookAhead system be in identifying adversarial contracts, can it produce a valid response quick enough so that defense actions could take place?
  3. 3)
    RQ3. How practical is it for individuals or organizations to deploy the LookAhead system and what does the performance of LookAhead look like in real-world scenarios?
6.2. RQ1: Effectiveness
6.2.1. Evaluation Metrics
In addition to commonly used binary classification performance metrics such as Accuracy, Precision, Recall and F1-Score, we also calculate False Positive Rates (FPR), which holds particular significance in our scenario.
6.2.2. Experiment Results
Table 1 presents our experiment results for evaluating the performance of LookAhead’s ML models. We calculate the performance metrics of various models used in the process of forming the LookAhead detection system, including $4$ candidate classifiers, a transformer-based classifier, and a meta classifier. Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest (RF) has the lowest recall ($0.7432$). Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across all metrics compared to other candidates. Its false positive rate reaches as low as $0.0041$, and it exhibits a high effective detection rate for adversarial samples. Other models also perform well. Except the one based on Decision Tree (DT), the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method. The meta classifier, K-nearest Neighbors Neural Network (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
6.2.3. Comparison with the Baseline
There have been studies that analyze smart contracts to detect contract vulnerabilities, but to the best of our knowledge, this is the first academic work that focus on analyzing smart contracts for the detection of adversarial contracts used by attackers across all types of DeFi attacks. The goal of both research directions is to prevent attackers from exploiting vulnerable contracts for profits, the only different is that our method aim to stop attacks from causing damage to potential targets while other works help potential targets reduce the possibility of being attacked. However, we emphasize that existing works on detecting vulnerabilities often focus on specific types of attacks/vulnerabilities, offering limited protection capabilities. Therefore, when comparing our results with the baseline, our goal is to: 1) Highlight the general applicability of LookAhead’s system in detecting all types of DeFi attacks, 2) Demonstrate the performance of LookAhead against previous state-of-art solutions.
Universality. We compare the performance of LookAhead in detecting all types of DeFi attacks with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts, and show that our system cannot not only match their efficacy in identifying reentrancy attacks, but can also offer similar effectiveness for detecting other types of attacks.
To evaluate LookAhead’s performance against reentrancy attack contracts, we tested our trained meta classifier model on all the reentrancy related adversarial contracts in our test dataset. We observe that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024).
Performance. The only similar research that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024). However, their approach relies solely on EVM bytecode while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts. We use Forta as the baseline for comparing the performance of our ML models.
We consider Forta an NLP-based method, hence we perform a direct comparison between their model and our transformer model first, thereafter we measure how much benefits can our meta classifier approach provide over Forta’s implementation. As the results in Table 1 indicate, our transformer alone can perform better and achieve higher F1-Socre than Forta, demonstrating the advantages of our Pruned Semantic-Control Flow Tokenization design.
Our meta classifier pushes the boundary even further by combining the capabilities of transformer model with the capabilities of other candidate classifiers to achieve F1-Score as high as $0.8966$. Furthermore, as described in §5.1, contract verification status feature introduces some bias into our dataset, to provide a fair comparison, we also evaluate the performance of our models with $verified$ feature removed. Experiments show that even with the feature $verified$ removed from our training and evaluation dataset, our classifiers still significantly outperform the most recent model published by Forta (released on February 6, 2023), achieving an F1-Socre of 0.8696, highlighting the superiority of LookAhead’s solution.
6.3. RQ2: Efficiency
Efficiency testing evaluates LookAhead’s resource consumption during its operating time, such as processing power requirements, memory usage, and detection throughput. All these contribute to delays in the identification of attackers and potential victims. Given the frequency of smart contract deployment transactions in decentralized finance, LookAhead should be able to quickly respond and react to potential attack activities within typical rescue time frames.
The most time-consuming step in LookAhead’s detection pipeline is the the contruction of PSCFT. Thanks to the rapid advancement of ML, classifier inference can be done on either CPUs or GPUs in a batched manner, which means in most cases it is a significantly faster process than binary lifting and analysis. Additionally, for external data augmentation pipelines that require sending API requests to outside sources, we can often easily employ multi-threading techniques to achieve high data retrieval concurrency. Thus, it can be assumed that the time required for analyzing a smart contract is roughly equivalent to the time needed for constructing its PSCFT.
To evaluate the efficiency, we measured the analysis time for all adversary contracts. The results showed that, except for one large contract with bytecode close to the maximum size, which took about $455$ seconds, most other contracts’ PSCFTs were constructed in under $20$ seconds, with an average time of $7.34$ seconds. Additionally, by considering the rescue time frames (cf. §3.2.4) of their corresponding DeFi attacks, we found that in $92.8\%$ of the cases, our system could respond in time by detecting the adversary contract before the adversary transaction occurred, leaving plenty of room for defense actions to be taken by affected parties and proving the viability of the LookAhead system.
6.4. RQ3: Practicality
We consider the practicality of the LookAhead system crucial to the validity of our research, therefore we conducted experiments measuring the detection performance of LookAhead to verify its effectiveness in realistic, dynamic environments.
From August 14 to August 18, 2024, we deployed the LookAhead system to a cloud server and conducted a live experiment to measure LookAhead’s performance in real-world scenarios. We use a meticulous review process that require us to 1) Wait for $2$ weeks to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, and we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks. These attacks caused over $50,000$ USD in financial losses. In terms of detection efficiency, except only one case where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction. Additionally, after thorough analysis, we observed that almost all the false positives are related to arbitrage, showing that the behaviour exhibited by arbitrage trading resembles some patterns in adversarial activities. For example, the dependency on flashloans, the high number of token ralated function calls, etc.
7. Threats to Validity
It should be noted that contracts may not necessarily need to be verified at the time of deployment, instead, contract creators might choose to complete the verification some time later. Limited by the API responses given by the blockchain explorers, we can only go as far as pinpointing the date of contract verification without knowing the exact time. Thus we cannot determine whether verification was done for a contract at the time of its deployment or not. However, based on our observations, most verified contracts have the same deployment and verification date, and even $6$ attackers verified their adversarial contracts, speculated to have occurred during deployment. Additionally, due to the scarcity of actual adversarial contract samples, the ratio between the two types of samples in our final dataset is approximately $1:40$. Furthermore, benign contracts were selected based on empirical observations and the assumption that adversarial contracts typically have fewer unique interactions with other accounts. However, there is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results.
8. Related Work
There have been many studies concerning the detection of abnormal contracts, primarily targeting Ponzi schemes and honeypots. They share similarities in methods which could be categorized into code similarity detection (Bartoletti et al., 2020; Sun et al., 2020), machine learning (Chen et al., 2020; Hu et al., 2022; Ji et al., 2023; Zheng et al., 2023), and symbolic execution (Chen et al., 2021). However, to the best of our knowledge, this is the first academic work that focuses on the detection of adversarial contracts across all types of DeFi attacks. Forta (For, 2024) also proposed a method that uses machine learning techniques to identify adversarial contracts, but their approach is limited to bytecode-level analysis. Aside from Forta, the only similar work is BlockWatchDog proposed by Yang et al. (Yang et al., 2024), which uses cross-contract static dataflow analysis technique to detect adversarial contracts in order to identify reentrancy vulnerabilities. However, their focus is solely on reentrancy attacks, which account for only $6.7\%$ of the attacks in our dataset, and their method relies heavily on knowing the victim addresses in advance.
9. Conclusion
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions. We build the first dataset for training PSCFT-based transformers, classifiers and meta classifiers for the binary classification of adversarial contracts. Based on this dataset, we found that the use of private mempool services by DeFi attackers is becoming increasingly common, rendering traditional transaction-based detection methods ineffective in most cases. Through the analysis of DeFi attack patterns, we extract a multi-dimensional feature set along with a special tokenization technique that is able to capture the intrinsic behaviours hidden in adversarial contracts. Utilizing these features, we build a transformer, multiple candidate classifiers and a meta classifier and evaluate them comprehensively. Experiments show that our method performs exceptionally well in detecting and preventing DeFi attacks with our models reaching F1-Scores as high as $0.8966$ and a false positive rate as low as $0.16\%$.

### v4

@@
 The ability to send private transactions was originally developed by researchers and industry practitioners to counter malicious activities on-chain such as front-running, but such services have unfortunately been increasingly abused by attackers for vulnerability exploitation purposes.
-Figure 1. The proportion of private transactions
-in DeFi attacks between 2022 and 2023.
+Figure 1. Percentage of DeFi attacks using
+private transactions between 2022 and 2023.
 Figure 2. Cumulative distribution function (CDF) of the rescue time frame in DeFi attacks.
@@
 We emphasize the critical importance of establishing a new approach for detecting DeFi attacks that can remain effective even when private mempool services are employed by attackers.
-Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor function during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims.
+Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims.
 As illustrated in Figure 2, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$.
@@
 Given that our dataset contains a combination of contracts from both recent and historical attacks, we believe that our feature selection will remain reasonably relevant and effective even in the rapid growing landscape of decentralized finance.
-The selection methods and criteria for building this dataset are described in details in section 4.
-These data have been made publicly available to facilitate future research in this domain.
 We build our classifiers using multiple supervised ML algorithms and conduct a comprehensive performance evaluation.
@@
   • A comprehensive dataset of adversarial and benign smart contracts with a variety of features pre-extracted, pre-processed and attached to each item.
-  • A custom designed smart contract bytecode lifting and analysis pipeline with integrated semantic tokenization protocol that is able to efficiently extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
+  • A smart contract bytecode lifting and analysis pipeline with an integrated protocol to extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
   • An ML-based system consisting of two open-weights classifiers and one meta classifier model that is pre-trained, fine-tunable, and capable of conducting accurate and just-in-time detection of potential adversarial attacks, even when private adversarial transactions are employed by attackers.
   • The implementation and a detailed evaluation as well as performance analysis of LookAhead’s detection system and methods.
-After providing some background in Section 2, we present an overview of LookAhead in Section 3.
-Section 4 details our dataset building strategies and Section 5 explains our methods, Section 6 reports our experimental results.
-After a brief discussion in Section 7 and 8, we conclude the paper in Section 9.
 2. Background
@@
 They are typically written in languages like Solidity (Dannen, 2017) and compiled into bytecode that can be executed within the EVM, a stack-based virtual machine supporting Turing-complete instructions.
+Tokens.
+Tokens are cryptocurrencies created using smart contracts.
+They come in two main types: fungible tokens and non-fungible tokens (NFT).
+Furthermore, Ether, used as the reward for anyone contributing in the consensus, is referred to as the native token.
+Decentralized Exchange (DEX).
+Unlike centralized exchanges (CEX), DEX doesn’t require users to deposit digital assets into the exchange for trading.
+Instead, it conducts asset transfers directly on the blockchain via smart contracts, enabling transparency, security, and decentralization.
 Flashloan.
@@
 The MEV-auction solution introduced by Flashbots (Fla, 2024) has recently gained popularity (Weintraub et al., 2022).
-This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others in the public mempool.
+This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others.
 However, such private mempool service are also utilized by attackers to send private adversarial transactions, aiming to execute attack logic and effectively bypass various detection methods.
@@
 3.1.2. Adversarial Patterns
-While DeFi attack logic can vary based on in the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives.
+While DeFi attack logic can vary based on the specific vulnerabilities being exploited, adversarial contracts often exhibit consistent behaviours due to the nature of their malicious intent and their profit motives.
 Our analysis of numerous adversarial contracts reveals a clustered set of patterns including the use of anonymous fund sources, closed-source code, and frequent token-related function calls.
@@
 Using the same method, we found 202 attacks (53.9%) with similar exposure behaviour in our dataset of $375$ adversarial contracts.
-Hence, we make one additional assumption about $\mathcal{A}$:
+Hence, we also consider a weaker adversary $\mathcal{A}^{\prime}$ satisfying one more assumption than $\mathcal{A}$:
   1. 2)
@@
 Our threat model leaves a short rescue time frame for detection and prevention to take place by the defenders and the targets, and enables the possibility of on-time reaction by victims of ongoing attacks.
-The rescue window is defined by $t^{window}=t_{first}-t_{deploy}-t^{pred}$, it has a typical value observed in our dataset of $t^{window}\geq 100$.
+The rescue window is defined by $t^{window}=t_{first}-t_{deploy}-t^{pred}$, it has a typical value observed in our dataset of $t^{window}\leq 100$.
 Within the rescue time frame several counterattack actions could be taken, we describe those in more details in §3.3.
@@
 After an ongoing attack has been identified by LookAhead, there are several defense techniques that can potentially be used by defenders and victim contract owners.
-1) If the victims’ addresses were found in the adversarial contract, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts.
+1) To defend against $\mathcal{A}^{\prime}$, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts.
 Moreover, if contract owners are willing to give us or any LookAhead system operator admin permissions to execute their shutdown methods on behalf of them, we can automatically react to possible attacks in an even more timely manner.
-2) We plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses.
+2) To counter $\mathcal{A}$, we plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses.
 If enough protocol users and developers incorporate our service as part of their access control pipeline, they can take full advantage of the real-time protection abilities of LookAhead at minimal to no cost to their operations.
@@
 First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and $selfdestruct\_count$.
-Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers) in combination with the PublicFunctionSelector relation, which maps public functions to their selectors, to determine the total number of public and private functions.
+Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers), with the help of PublicFunctionSelector, which maps public functions to their selectors, to determine the total number of public and private functions.
 This combination also helps identify flashloan callback functions.
@@
 Based on the IR of a contract, we construct a function-by-function visualization $F=\{F^{private}\cup F^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements.
-Functions have 4-byte signatures derived from the Keccak hash (SHA3) of their canonical names and can be either private or public.
+Functions have 4-byte selector derived from the Keccak hash (SHA3) of their canonical names and can be either private or public.
 Each function contains multiple basic blocks (BB) chained together ($\{\mathcal{F}_{i}=\{BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|\mathcal{F}_{i}|}\}\}$), and each block contains multiple statements.
@@
 The pruning process starts with the sorting and reordering of functions based on function types (public first, then private) and their canonical names.
-Gigahorse can only recover the names for some of the public functions and external call target methods, hence if the derivation of canonical names from function signatures failed for some public function $\mathcal{F}^{public}_{i}\in F^{public}$, then $\mathcal{F}^{public}_{i}$ will be given a name of $UnknownFunction_{i}$.
+Gigahorse can only recover the names for some of the public functions and external call methods, hence if the derivation of canonical names from function selectors failed, we will leave them as-is and recover their names in the semantic recovery stage.
 For any private function $\mathcal{F}^{private}_{j}\in F^{private}$, we give it a canonical name of $InternalFunction_{i}$.
@@
 For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label.
-Contract developers might choose to store the target address in the EVM storage or as a variable/constant which affects our ability to directly perform address matching.
-We use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage.
+Contract developers might choose to store the target address in the EVM storage or as a variable/constant, in which case we use the same technique as (Kong et al., 2023) to run a Datalog program that finds the slot location storing the address value and retrieves the value from the target storage.
 If this process did not result in a matching label, we assign the address a default label UnknownTarget.
-Function signatures are generated by hashing the function names with a secure hashing algorithm, and the hashing process is a one-way trap door.
+Function selectors are generated by hashing the function signatures with a secure hashing algorithm, and the hashing process is a one-way trap door.
 However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings.
-Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their Application Programming Interfaces (API) upon failed recovery by Gigahorse.
+Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their APIs upon failed recovery by Gigahorse.
+If the recovery still fails, we fallback to the default label UnknownFunc.
 If the aforementioned process was completed successfully, the external call statement will be converted to Balancer.flashLoan(…args).
@@
 By feeding the predictions produced by the two base models into the meta model, we are able to 1) Combine the strengths of each model and automatically learn the optimal way to construct the inference result based on model significance, 2) Optimize the final classification boundaries to achieve more accurate detection of adversarial contracts.
-We use K-Nearest Neighbors (KNN) as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space.
+We use K-Nearest Neighbors (KNN) to as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space.
 KNN is well-suited for combining the outputs of different classifiers such as the transformer model and one of the candidate classifiers, because it is able to combine the complementary strengths of both models by analyzing their predictions and finding the optimal decision boundary.
@@
 Consequently, traditional cross-validation techniques, which assume input values are independent with respect to each other, are not suitable for our evaluation.
-Instead, to achieve such an objective, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing.
+Instead, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing.
 Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
@@
 Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest (RF) has the lowest recall ($0.7432$).
-Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across all metrics compared to other candidates.
+Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across most metrics compared to other candidates.
 Its false positive rate reaches as low as $0.0041$, and it exhibits a high effective detection rate for adversarial samples.
@@
 Except the one based on Decision Tree (DT), the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method.
-The meta classifier, K-nearest Neighbors Neural Network (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
+The meta classifier, K-nearest Neighbors (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
 6.2.3. Comparison with the Baseline
@@
 Universality.
-We compare the performance of LookAhead in detecting all types of DeFi attacks with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts, and show that our system cannot not only match their efficacy in identifying reentrancy attacks, but can also offer similar effectiveness for detecting other types of attacks.
-To evaluate LookAhead’s performance against reentrancy attack contracts, we tested our trained meta classifier model on all the reentrancy related adversarial contracts in our test dataset.
-We observe that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024).
+We compare the performance of LookAhead with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts to demonstrate the universality of the LookAhead system.
+When our trained meta classifier model was tested on all the reentrancy related adversarial contracts in our test dataset, we observed that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024).
+This observation along with the experiment results presented in §6.2.2 show that despite being trained on contracts used in a variety of attacks of different types (cf.
+§4.1.1), LookAhead still achieves comparable performance in detecting any specific types of DeFi attacks.
 Performance.
-The only similar research that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024).
+The only similar work that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024).
 However, their approach relies solely on EVM bytecode while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts.
@@
 From August 14 to August 18, 2024, we deployed the LookAhead system to a cloud server and conducted a live experiment to measure LookAhead’s performance in real-world scenarios.
-We use a meticulous review process that require us to 1) Wait for $2$ weeks to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
-Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, and we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks.
+We use a meticulous review process that require us to 1) Wait for $2$ weeks (September 1) to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
+Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, in which we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks.
 These attacks caused over $50,000$ USD in financial losses.
@@
 For example, the dependency on flashloans, the high number of token ralated function calls, etc.
-7. Threats to Validity
-It should be noted that contracts may not necessarily need to be verified at the time of deployment, instead, contract creators might choose to complete the verification some time later.
+7. Discussion
+Threats to Validity.
 Limited by the API responses given by the blockchain explorers, we can only go as far as pinpointing the date of contract verification without knowing the exact time.
-Thus we cannot determine whether verification was done for a contract at the time of its deployment or not.
-However, based on our observations, most verified contracts have the same deployment and verification date, and even $6$ attackers verified their adversarial contracts, speculated to have occurred during deployment.
+Thus we cannot determine whether verification was done at the time of deployment or not.
 Additionally, due to the scarcity of actual adversarial contract samples, the ratio between the two types of samples in our final dataset is approximately $1:40$.
+This imbalance could affect the performance of our models.
 Furthermore, benign contracts were selected based on empirical observations and the assumption that adversarial contracts typically have fewer unique interactions with other accounts.
-However, there is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results.
+There is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results.
+Smart contracts can be created through internal transactions triggered during the execution of a normal transaction.
+Those contracts are not included in our dataset.
+Limitations.
+Our current method has certain limitations in terms of the spectrum of adversarial contracts that LookAhead is capable of detecting.
+When selecting features, we rely on existing attack patterns seen in DeFi attacks involving financial losses and those in prevalent attack scenarios.
+Consequently, there is a probability that we might have missed attacks that do not result in financial damages and certain adversarial contracts with novel attack vectors.
+Furthermore, in the cat-and-mouse game of DeFi attack and defense, attackers may evade our detection by using techniques like code obfuscation to alter the CFG of contract programs or by utilizing KYC-required exchanges as funding sources that makes it harder for our system to recognize adversarial intentions.
 8. Related Work
@@
 They share similarities in methods which could be categorized into code similarity detection (Bartoletti et al., 2020; Sun et al., 2020), machine learning (Chen et al., 2020; Hu et al., 2022; Ji et al., 2023; Zheng et al., 2023), and symbolic execution (Chen et al., 2021).
-However, to the best of our knowledge, this is the first academic work that focuses on the detection of adversarial contracts across all types of DeFi attacks.
+However, to the best of our knowledge, this is the first academic work that focuses on identifying adversarial contracts, and the first to support detection across all types of DeFi attacks.
 Forta (For, 2024) also proposed a method that uses machine learning techniques to identify adversarial contracts, but their approach is limited to bytecode-level analysis.

### v6

@@
 Abstract.
-Decentralized Finance (DeFi) incidents stemming from the exploitation of smart contract vulnerabilities have culminated in financial damages exceeding $3$ billion US dollars.
-Existing defense mechanisms typically focus on detecting and reacting to malicious transactions executed by attackers that target victim contracts.
+The exploitation of smart contract vulnerabilities in Decentralized Finance (DeFi) has resulted in financial losses exceeding $3$ billion US dollars.
+Existing defense mechanisms primarily focus on detecting and reacting to adversarial transactions executed by attackers that target victim contracts.
 However, with the emergence of private transaction pools where transactions are sent directly to miners without first appearing in public mempools, current detection tools face significant challenges in identifying attack activities effectively.
-Based on the fact that most attack logic rely on deploying one or more intermediate smart contracts as supporting components to the exploitation of victim contracts, in this paper, we propose a new direction for detecting DeFi attacks that focuses on identifying adversarial contracts instead of adversarial transactions.
-Our approach allows us to leverage common attack patterns, code semantics and intrinsic characteristics found in malicious smart contracts to build the LookAhead system based on Machine Learning (ML) classifiers and a transformer model that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions of potential zero-day attacks.
-Our contributions are three-fold: First, we construct a comprehensive dataset consisting of features extracted and constructed from $210,643$ recent contracts and $375$ hand-picked then carefully reviewed and identified adversarial contracts deployed on the Ethereum and Binance Smart Chain (BSC) blockchains.
-Secondly, we design a condensed representation of smart contract programs called Pruned Semantic-Control Flow Tokenization (PSCFT) and use it to train a combination of ML models that understand the behaviour of malicious codes based on function calls, control flows and other pattern-conforming features.
-Lastly, we provide the complete implementation of LookAhead and the evaluation of its performance metrics for detecting adversarial contracts.
-Our experiments show that LookAhead achieves an F1-score as high as $0.8966$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta, with a False Positive Rate (FPR) at only $0.16\%$.
-copyright: none
+Based on the fact that most attack logic rely on deploying intermediate smart contracts as supporting components to the exploitation of victim contracts, novel detection methods have been proposed that focus on identifying these adversarial contracts instead of adversarial transactions.
+However, previous state-of-the-art approaches in this direction have failed to produce results satisfactory enough for real-world deployment.
+In this paper, we propose LookAhead, a new framework for detecting DeFi attacks via unveiling adversarial contracts.
+LookAhead leverages common attack patterns, code semantics and intrinsic characteristics found in adversarial smart contracts to train Machine Learning (ML)-based classifiers that can effectively distinguish adversarial contracts from benign ones and make timely predictions of different types of potential attacks.
+Experiments on our labeled datasets show that LookAhead achieves an F1-score as high as $0.8966$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution, with a False Positive Rate (FPR) at only $0.16\%$.
+Blockchain, DeFi, Smart contract, Malware detection
+*Jian Liu is the corresponding author.
+copyright: ccdoi: 10.1145/3729353journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE083publicationmonth: 7ccs: Security and privacy Intrusion detection systems
 1. Introduction
-Decentralized Finance (DeFi) has gained significant traction within the blockchain ecosystem over the recent years and is continuing to evolve rapidly, stimulating the rise of a diverse array of DeFi applications such as: decentralized exchanges (DEXs) (Adams et al., [n.
-d.]; Martinelli and Mushegian, [n.
-d.]), lending platforms (AAV, 2024; Robert and Geoffrey, [n.
-d.]), and derivatives instruments (Juliano, [n.
-d.]).
+Decentralized Finance (DeFi) has gained significant traction within the blockchain ecosystem over the recent years, stimulating the rise of a diverse array of DeFi applications such as: decentralized exchanges (DEXs) (Adams et al., 2021; Martinelli and Mushegian, 2019) and lending platforms (AAV, 2020; Robert and Geoffrey, 2019).
 The total value locked (TVL) in these DeFi protocols has exceeded \$100 billion in mid 2024 (Def, 2024).
-Among the various blockchain networks, Ethereum (Wood et al., 2014) and Ethereum Virtual Machine (EVM)-compatible chains (Arb, 2024; BNB, 2024; Opt, 2024; Pol, 2024) have formed a particularly thriving DeFi ecosystem, collectively they play a pivotal role in shaping the future of the blockchain landscape.
+Among the various blockchains, Ethereum (Wood et al., 2014) and Ethereum Virtual Machine (EVM)-compatible chains (Arb, 2024; BNB, 2024; Opt, 2024; Pol, 2024) have formed a particularly thriving DeFi ecosystem.
 At the heart of Decentralized Finance are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries.
@@
 Targeting smart contracts, there are mainly two types of attack patterns commonly used by DeFi attackers: an attacker can either directly invoke specific public functions of the victim contracts in order to exploit a vulnerability (Tre, 2024; Sus, 2024b), or they can deploy an adversarial contract containing the complete attack logic onto the blockchain first, then initiate the attack through the intermediate contract by calling their entry functions afterwards (Bea, 2024; Bur, 2024; Pan, 2024; Sus, 2024a).
-Limitations of Existing Work.
-Existing real-time defense mechanisms leverage heuristics (Qin et al., 2023b; Wu et al., 2023; Zhang et al., 2020; Zhou et al., 2020) and Machine Learning (ML) (Gai et al., 2023; Wang et al., 2022) techniques to detect and respond to yet-to-be-confirmed transactions in public mempools that are adversarial in nature.
-The main issue with this transaction-based detection approach is that it faces significant challenges in identifying private adversarial transactions.
-Namely, attackers can utilize private mempool services to send adversarial transactions directly to miners, evading visibility from other participants of the blockchain network before the transaction is confirmed and effectively bypassing attack detection.
-The ability to send private transactions was originally developed by researchers and industry practitioners to counter malicious activities on-chain such as front-running, but such services have unfortunately been increasingly abused by attackers for vulnerability exploitation purposes.
-Figure 1. Percentage of DeFi attacks using
-private transactions between 2022 and 2023.
-Figure 2. Cumulative distribution function (CDF) of the rescue time frame in DeFi attacks.
-Key Observations.
+Detection Based on Adversarial Transactions.
+Most existing real-time defense mechanisms leverage heuristics (Qin et al., 2023b; Wu et al., 2023; Zhang et al., 2020; Zhou et al., 2020) and Machine Learning (ML) (Gai et al., 2023; Wang et al., 2022) techniques to detect and respond (e.g.
+via front-running) to yet-to-be-confirmed attack transactions in public mempools.
+The main issue with this transaction-based detection approach is that it faces significant challenges in applying to private adversarial transactions.
+Namely, attackers can utilize private mempool services to send transactions directly to miners, evading visibility from other participants of the blockchain network before it’s confirmed and effectively bypassing attack detection and avoid any preventive actions from taking place.
 Based on our empirical study of the historical DeFi incidents over the past few years, we found that out of the $142$ DeFi attacks on Ethereum collected by us, $80$ cases ($56.3\%$) involve transactions using private mempool services.
-Moreover, as shown in Figure 2, the proportion of adversarial transactions sent using private services exhibits a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023.
-We emphasize the critical importance of establishing a new approach for detecting DeFi attacks that can remain effective even when private mempool services are employed by attackers.
-Another important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor during adversarial contract deployment, which provides a detection and rescue time frame for defenders and potential victims.
-As illustrated in Figure 2, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$.
-Based on the aforementioned studies and observations, we design our method for preventing DeFi attacks via unveiling adversarial contracts.
+Moreover, the proportion of adversarial transactions sent using private services is observed to exhibit a significant upward trend, rising from $44.4\%$ in the first half of 2022 to a noticeably higher $77.8\%$ in the latter half of 2023.
+Detection Based on Adversarial Contracts.
+The ability for an attacker to execute private adversarial transactions severely undermines the detection capabilities of existing transaction-based solutions, rendering them largely ineffective.
+In contrast, some works have redirected their focus towards developing detection strategies based on identifying adversarial contracts.
+As an example, Forta (For, 2024) applies NLP techniques and uses a simple logistic regression model to analyze contract bytecode for malicious intents, but it fails to achieve satisfactory performance (Table 1).
+Beyond Forta, the only comparable work is BlockWatchDog (Yang et al., 2024), which only support detecting reentrancy attacks.
+We emphasize the critical importance of designing an effective solution for detecting a wide range of DeFi attacks that can remain functional even when private mempool services are employed by the attackers.
 Our method.
-In this paper, we propose the LookAhead system, representing a new direction for detecting DeFi attacks, i.e., focusing on identifying adversarial contracts instead of adversarial transactions, and allowing us to proactively identify potential attack intentions and victims, even in the case of private adversarial transactions.
-By shifting the focus to analyzing and identifying adversarial contracts, we filled several gaps currently present in traditional transaction-based detection tools.
-We first make an empirical observation that adversarial contracts used in DeFi attacks often contain similar patterns, for example, over 70% of the contracts in our adversarial contract dataset use funds from anonymous sources for deployment; more than 98% are closed-sourced, etc.
-We also note that depending on the vulnerabilities targeted, attackers write contract code in specific manners.
-For instance, flashloan and reentrancy attacks are two of the most notorious and frequently studied exploitation.
-Both of them require the establishment of complex function call chains and the implementation of attack logic within callback functions.
-To use these behaviours to our advantage for identifying adversarial contracts, we using the stacking technique to design three ML models, one transformer, multiple candidate classifiers, and a meta classifier that, when used together, can understand the intrinsic characteristics of malicious smart contracts based on code semantics and common patterns.
-To support the ML models in understanding the behaviour of Solidity smart contract code, we develop a specialized tokenization method called Pruned Semantic-Control Flow Tokenization (PSCFT) that is integrated directly into a custom contract bytecode decompiler.
-By leveraging modern ML architecture and models, we can produce batch predictions for newly created contracts on-chain in a timely manner with amortized $t^{pred}\ll 100s$.
-The result of our design is LookAhead, a system that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions of potential zero-day attacks.
-To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our attack contract dataset.
+In this paper, we propose the LookAhead system, a new framework for detecting DeFi attacks via unveiling adversarial contracts.
+It achieves high detection effectiveness through systematic feature design and selection, the construction of a comprehensive contract dataset, and a combination of advanced classifier models.
+In §3.2, we describe in detail the threat model assumptions and the scope of detection followed by our proposed system.
+We make an empirical observation that adversarial contracts used in DeFi attacks often contain similar patterns ($>70\%$ of the adversarial contracts use funds from anonymous sources for deployment; $>98\%$ are closed-sourced, etc.) We also note that depending on the vulnerabilities targeted, attackers write contract code in specific manners.
+For instance, flashloan attacks rely on complex call chains and the implementation of attack logic within callback functions.
+To use these behaviours to our advantage for identifying adversarial contracts, we using the stacking technique to design three ML models (transformer, classifier, and a meta classifier) that, when used together, can understand the intrinsic characteristics of malicious contracts based on code semantics and common patterns.
+To support the models in understanding the behaviour of Solidity code, we develop a specialized tokenization method called Pruned Semantic-Control Flow Tokenization (PSCFT) that is integrated directly into a contract bytecode decompiler.
+An important observation made by (Zhou et al., 2023) is that most attacks are not executed atomically within the constructor during adversarial contract deployment, which provides a rescue time frame (after an adversarial contract has been deployed and before the attacker execute the attack logic via the contract) for defenders and victims.
+We use $tx_{deploy}$ to denote a contract deployment transaction, and $tx_{first}$ to represent the transaction in which the attacker initiates the actual attack via the adversarial contract.
+In our dataset, over $60\%$ of attacks exhibit a difference between the confirmation time of $tx_{deploy}$ and $tx_{first}$ that satisfies: $t_{first}-t_{deploy}\geq 100s$.
+By designing an efficient feature extraction pipeline and leveraging modern ML architectures, we can produce predictions for newly created contracts on-chain in a timely manner with amortized $t^{pred}\ll 100s$.
+The result of our design is LookAhead, a system that is able to effectively distinguish adversarial contracts from benign ones, and make just-in-time predictions ($t_{deploy}+t^{pred}<t_{first}$) of potential attacks.
+To ensure the performance and generality of our ML models, we hand-picked then carefully reviewed and identified $375$ adversarial contracts used in DeFi incidents between April 2020 and July 2024 from both the Ethereum and Binance Smart Chain (BSC) to form our adversarial contract dataset.
 And in total $796,437$ contracts deployed on Ethereum from June 2022 to June 2024 were collected, with $210,643$ of them being used as benign samples after filtration.
@@
 We build our classifiers using multiple supervised ML algorithms and conduct a comprehensive performance evaluation.
-Experiment results show that our method for identifying adversarial contracts performs exceptionally well with an F1-Score produced by the KNN meta classifier reaching as high as $0.8966$, and a false positive rate of only $0.16\%$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta.
+Experiment results on our labeled datasets show that our method for identifying adversarial contracts performs exceptionally well with an F1-Score produced by the KNN meta classifier reaching as high as $0.8966$, and a false positive rate of only $0.16\%$, which represents an improvement of over $44.4\%$ compared to the previous state-of-the-art solution Forta.
 Contributions. We summarize our contributions as follows:
-  • We propose a new direction for detecting DeFi attacks that focuses on identifying adversarial contracts instead of adversarial transactions.
-  • A comprehensive dataset of adversarial and benign smart contracts with a variety of features pre-extracted, pre-processed and attached to each item.
-  • A smart contract bytecode lifting and analysis pipeline with an integrated protocol to extract and generate features and PSCFT for transformer and classifier training and fine-tuning.
-  • An ML-based system consisting of two open-weights classifiers and one meta classifier model that is pre-trained, fine-tunable, and capable of conducting accurate and just-in-time detection of potential adversarial attacks, even when private adversarial transactions are employed by attackers.
-  • The implementation and a detailed evaluation as well as performance analysis of LookAhead’s detection system and methods.
+  • We highlight the limitations in existing transaction and contract-based detection methods and propose a new framework LookAhead that achieves significantly better performance;
+  • We design a smart contract bytecode lifting and analysis pipeline with an integrated protocol to extract features and generate PSCFT for Natural Language Processing (NLP) training;
+  • We construct the first large-scale comprehensive dataset of adversarial (manually-labelled) and benign (methodically selected) smart contracts consisting of a set of useful contract features based on extensive empirical observations with statistical data support;
+  • We build and evaluate an ML-based system consisting of two open-weights classifiers and one meta classifier model, the results are compared with previous state-of-the-art works to substantiate our system’s effectiveness and practicality.
 2. Background
 2.1. Ethereum & Decentralized Finance (DeFi)
-Blockchain is a distributed ledger built on top of a peer-to-peer (P2P) network, maintained through consensus mechanisms to ensure the consistency of the ledger.
+Blockchain is a decentralized ledger maintained over a peer-to-peer network via consensus mechanisms.
 Ethereum (Wood et al., 2014) is one of the most used blockchain platforms.
@@
 DeFi is a blockchain-based financial ecosystem powered by smart contracts to offer financial services in a more open and transparent manner.
-There are several unique concepts in DeFi that differ from traditional finance.
 Accounts.
@@
 Transactions are used to transfer Ether, invoke contracts, or create contracts.
-After a transaction is sent, it is first placed in the mempool of a blockchain node.
-Public mempools are transparent and openly accessible and filled with pending transactions, waiting to be selected and finalized by block producers.
+After a transaction is sent, it is first placed in the mempool of a blockchain node, waiting to be selected and finalized by block producers.
 Smart Contracts.
@@
 2.2. Maximal Extractable Value (MEV)
-Traditionally, miner determine the order of transactions in a block based on gas prices.
-Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the gas price, allowing for front-running (Eskandari et al., 2020; Torres et al., 2021) or back-running to happen.
-The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Maximal Extractable Value (MEV) (Daian et al., 2020).
-It has been proven to have a significant impact on the security of blockchain networks (Qin et al., 2022).
+Traditionally, miners determine the order of transactions in a block based on gas prices.
+Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the gas price, allowing for front-running (Eskandari et al., 2020; Torres et al., 2021).
+The profit gained from such manipulation, beyond standard rewards and fees, is known as Maximal Extractable Value (Daian et al., 2020).
 Front-running.
-Front-running is a fundamental means used to extract MEV, typically employed in arbitrage, liquidation, and sandwich attacks.
-Front-runners monitor the public mempool for victim transactions that meets specific conditions, and attempt to raise the gas price to execute their own transactions first against the original transaction.
-Furthermore, front-running has also been used in intrusion prevention systems (Qin et al., 2023a; Zhang et al., 2023), as a way to front-run adversarial transactions in order to prevent DeFi attacks.
+Front-running is a fundamental means used to extract MEV, where bots monitor the public mempool for target transactions and raise gas prices to execute their own first.
+It has also been used in transaction-based intrusion prevention systems (Qin et al., 2023a; Zhang et al., 2023) to front-run adversarial transactions, thereby preventing DeFi attacks.
+Figure 1. Read-only reentrancy attack on Conic Finance, using a private adversarial transaction to evade detection, could have been prevented during the rescue time frame.
 Private Mempool Services.
-In response to MEV, researchers and industry practitioners have proposed various countermeasures (Yang et al., 2022).
-The MEV-auction solution introduced by Flashbots (Fla, 2024) has recently gained popularity (Weintraub et al., 2022).
-This solution introduced the concept of relay systems (Lyu et al., 2022), which provide private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others.
-However, such private mempool service are also utilized by attackers to send private adversarial transactions, aiming to execute attack logic and effectively bypass various detection methods.
+In response to MEV, Flashbots’ MEV-auction has been widely adopted (Fla, 2024; Weintraub et al., 2022).
+This solution provides private channels for transactions to be submitted directly from the users to block producers without being broadcasted to the network, ensuring that transactions remain private until they are finalized, preventing them from being monitored or intercepted by others.
+However, these services have increasingly been abused by attackers to conceal adversarial transactions, effectively evading mainstream transaction-based detection methods.
 3. Overview of LookAhead
 3.1. Motivating Example
-Figure 3. Read-only reentrancy attack on Conic Finance, using a private adversarial transaction to evade detection, could have been prevented during the rescue time frame.
 On July 21, 2023, DeFi protocol Conic Finance suffered a major exploit (DLNews, 2024) causing financial losses exceeding \$3 million.
@@
 Notably, the adversary transaction that triggered the attack was labeled as an MEV Transaction on Etherscan  [^1] https://etherscan.io/tx/0x8b74995d1d61d3d7547575649136b8765acb22882960f0636941c44ec7bbe146, indicating the use of private mempool services.
-This allowed the attacker to bypass the public mempool and invalidate any potential front-running defense attempts.
-This attack highlights the limitations of traditional attack prevention mechanisms that rely on detecting adversarial transactions, particularly when private transactions are used.
+This allowed the attacker to bypass the public mempool and invalidate any potential front-running defense attempts, highlighting the limitations of transaction-based prevention mechanisms that rely on detecting adversarial transactions.
 3.1.1. Rescue Possibility
-Despite the attacker using private mempool services to conceal the adversarial transaction, there was still an opportunity to prevent the attack, based on the following facts: 1) The ConicEthPool contract’s address — the vulnerable target — was written in plaintext in the adversary contract at the time of contract deployment; 2) There was a 36-second window (3 blocks) between the deployment of the adversarial contract and the execution of the adversarial transaction; 3) The ConicEthPool contract had an access-controlled shutdownPool function, which offers the ability to perform an emergency pause to block the adversary’s attempt to call the deposit function.
+Despite the attacker using private mempool services to conceal the adversarial transaction, there was still an opportunity to prevent the attack, based on the following facts: 1) The ConicEthPool contract’s address — the vulnerable target — was written in plaintext in the adversarial contract at deployment; 2) There was a 36-second window between the deployment of the adversarial contract and the execution of the adversarial transaction; 3) The ConicEthPool contract had a shutdownPool function, which offers the ability to perform an emergency pause to block the adversary’s attempt to call the deposit function.
 Thus, detecting the adversarial contract at the deployment stage, before the confirmation of the adversarial transaction, could have prevented the exploit, regardless of whether the attacker used a private mempool service or not.
@@
 Additionally, adversarial contracts often have distinctive interaction logic with victim contracts, which differs significantly from the patterns observed in normal contracts, resulting in easily distinguishable code structures and external call flows.
-Based on these observations, ML techniques can be employed to effectively identify adversarial contracts — regardless of the specific attack type — by learning and predicting these common patterns.
+Based on these observations, ML techniques can be employed to effectively identify adversarial contracts (regardless of the specific attack type) by learning and predicting these common patterns.
 3.2. Threat Model
 3.2.1. DeFi Attacks
-There are many types of security breaches with complicated root causes that could lure attackers into exploiting a victim contract.
-In different attack scenarios, an adversary could use different techniques to conduct their exploitation.
+There are many types of security breaches with complicated root causes, potentially leading to attacks on victims.
+In various scenarios, an adversary could use distinct techniques for exploitation.
 In this section, we present these attack methods, define key concepts involved in the attack process, and restrict LookAhead’s detection scope to a certain area of focus.
@@
 This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failures thanks to the atomic nature of a blockchain transaction.
-Using this method, an attacker first deploys a smart contract on-chain containing the core attack logic, then initiate the attack by calling the entry point function of the adversarial contract.
+Using this method, an attacker first deploys a smart contract on-chain containing the core attack logic, then initiate the adversarial transaction by calling the entry point function of the adversarial contract.
 Most DeFi attacks follow this pattern, hence we focus on addressing this type of attacks.
 Scope of Detection.
-We specify three attack scenarios that fall outside of the scope this study.
+We specify two attack scenarios that fall outside of the scope this study.
 1) Attacks that do not require contract deployments.
@@
 Since the adversarial contracts used in these attacks already exhibit execution traces of the actual attack at the time of deployment, we consider them as adversarial transactions, which fall outside the scope of our detection.
-3) We do not place our attention on detecting attacks such as rug pulls as they can only be launched by privileged users.
-LookAhead focuses on unveiling attacks with intention to cause financial loss for DeFi protocols by exploiting vulnerabilities in on-chain smart contracts.
-To identify an attack, our goal is to detect adversarial contracts that contain the core attack logic.
-Other contracts, such as those used solely for concealing profits to avoid detection, fake token contracts to facilitate the main attack, and counterattack contracts used by front-running bots will not be included in our analysis.
+LookAhead focuses on unveiling attacks that intend to cause financial loss for DeFi protocols by exploiting vulnerabilities in on-chain smart contracts.
+To identify an attack, our solution detects adversarial contracts that contain the core attack logic.
+Other contracts, such as those used solely for concealing profits, fake token contracts that facilitate the main attack, and counterattack contracts used by front-running bots will not be included for our analysis.
 3.2.2. Adversaries
 We consider a resource bounded adversary $\mathcal{A}$ that is financially rational, capable of deploying an arbitrary amount of smart contracts and execute both public and private transactions on EVM-compatible blockchains.
-Attacker Assumptions. Based on previous research findings (Zhou et al., 2023), our study makes the following assumption about $\mathcal{A}$:
+Attacker Assumptions. Based on previous research findings (Zhou et al., 2023) (IEEE S&P’23), our study makes the following assumption about $\mathcal{A}$:
   1. 1)
-    $\mathcal{A}$ do not initiate attacks by batching $tx_{deploy}$ and $tx_{first}$ together in an atomic transaction during contract deployment, meaning that $\mathcal{A}$ splits an attack into two stages following the second pattern described in Figure 4;
+    $\mathcal{A}$ does not initiate attacks by batching $tx_{deploy}$ and $tx_{first}$ together in an atomic transaction during contract deployment, meaning that $\mathcal{A}$ splits an attack into two stages following the second pattern described in Figure 2;
 In a prior research (Parhizkari et al., 2023), it was observed that 79.49% of attacks in their dataset exposed the victim contract address before the adversarial transaction.
@@
   1. 2)
-    $\mathcal{A}$ hard-code exploitation targets’ addresses in their malicious contracts whenever possible, and the victim’s address will emerge before the first attack transaction $tx_{first}$.
+    $\mathcal{A}^{\prime}$ hard-codes exploitation targets’ addresses in their adversarial contracts whenever possible, and the victim’s address will emerge before the first adversarial transaction $tx_{first}$.
 3.2.3. Victim of Attacks
@@
 We do not use any source code information from the contracts, but only the bytecode deployed on the blockchain to ensure our model is capable of analyzing any smart contract on EVM blockchains.
-To do that, LookAhead uses the open-source binary lifting and analysis library Gigahorse (Grech et al., 2019) to decompile the contract bytecode and generate intermeidate representation (IR).
-We then run a program to visualize the Control Flow Graph (CFG) based on the IR.
+To do that, LookAhead uses the binary lifting and analysis library Gigahorse (Grech et al., 2019) to decompile the contract bytecode and generate intermeidate representation (IR).
 PSCFT and Feature Extraction.
-With the basic CFG, we use a specialized feature extraction program to pull out features including but not limited to: token-related call information, flashloan usage information, etc.
-Finally, we convert the CFG into a condensed text form (PSCFT) and populate it with semantic information such as function names and external call target address labels.
+With the raw IR, we use a specialized feature extraction program to pull out features including but not limited to: token-related call information, flashloan usage information, etc.
+Finally, we convert the IR into a condensed text form (PSCFT) and populate it with semantic information such as function names and external call target address labels.
 Classifier Inference.
@@
 These ML models are made up of a transformer, an optimal candidate classifier, and a meta classifier trained using supervised machine learning methods for binary classification.
-They are able to output either the classification result $label_{pred}\in\{0,1\}$ indicating whether the contract is adversary ($1$) or not (00), or they can produce a confidence score $0\%\leq p_{pred}\leq 100\%$ representing the possibility of the input contract being malicious.
+They are able to output either the classification result $label_{pred}\in\{0,1\}$ indicating whether the contract is adversarial ($1$) or not (00), or they can produce a confidence score $0\%\leq p_{pred}\leq 100\%$ representing the possibility of the input contract being adversarial.
 Defense Mechanisms.
-After an ongoing attack has been identified by LookAhead, there are several defense techniques that can potentially be used by defenders and victim contract owners.
-1) To defend against $\mathcal{A}^{\prime}$, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts.
-Moreover, if contract owners are willing to give us or any LookAhead system operator admin permissions to execute their shutdown methods on behalf of them, we can automatically react to possible attacks in an even more timely manner.
-2) To counter $\mathcal{A}$, we plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses.
-If enough protocol users and developers incorporate our service as part of their access control pipeline, they can take full advantage of the real-time protection abilities of LookAhead at minimal to no cost to their operations.
+After detecting an ongoing attack, LookAhead enables several potential defense mechanisms for defenders and victim contract owners: 1) To defend against $\mathcal{A}^{\prime}$, notifications are sent to targets, allowing them to take defensive actions, such as executing an emergency pause.
+If contract owners grant admin permissions to LookAhead, the system can autonomously trigger shutdown methods on behalf of them.
+2) To mitigate against $\mathcal{A}$, we propose a blacklist service for DeFi protocols, enabling them to block transactions from flagged adversarial contracts.
 4. Dataset Building
 4.1. Contract Collection
-To perform supervised training of our classifier models, we collect a large set of past and recent smart contracts.
-For the positive (adversarial) dataset, we curated a collection of adversarial contracts used in past real-world DeFi incidents.
-For the negative (benign) dataset, we follow a methodical process to choose samples that are very unlikely to be related to any attacks.
+To perform supervised training of our classifier models, we collect a large set of smart contracts from the Ethereum and BSC.
+For the positive (adversarial) dataset, we curated a collection of adversarial contracts used in real-world DeFi incidents.
+For the negative (benign) dataset, we follow a methodical process to collect contract samples that are very unlikely to be related to any attacks.
 4.1.1. Adversarial Contracts
-To collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as GitHub repositories documenting DeFi incidents (DeF, 2024; Lea, 2024), and alerts from security companies (blo, 2024; cer, 2024; SlowMist, [n.
-d.]).
-We conduct a meticulous review of these incidents to identify adversarial contracts used by attackers.
+To collect adversarial contracts, we gather information on DeFi incidents from various publicly available sources, such as GitHub repositories documenting DeFi incidents (DeF, 2024; Lea, 2024), and alerts from security companies (blo, 2024; cer, 2024; slo, 2024).
+We conduct a meticulous review of these incidents to identify the contracts used by attackers.
 Between April 2020 and July 2024, we collected 506 DeFi attacks.
-Of these, $29$ attacks ($5.7\%$) involved direct attacks from EOAs without using contracts, and $37$ attacks ($7.3\%$) were executed directly within the contract constructor.
+In these, $29$ attacks ($5.7\%$) involved direct attacks from EOAs without using contracts, and $37$ attacks ($7.3\%$) were executed directly within the contract constructor.
 Both cases fall outside the scope of our detection.
@@
 This left us with $375$ adversarial contracts: $142$ from Ethereum and $233$ from BSC.
-Following prior research (Zhou et al., 2023), we classify the attacks into three categories: Untrusted or unsafe calls ($13.3\%$), Access control mistakes ($14.4\%$), and Coding mistake ($72.3\%$).
+Following prior research (Zhou et al., 2023), we classify the attacks into three categories: 1) Untrusted or unsafe calls ($13.3\%$), 2) Access control mistakes ($14.4\%$), and 3) Coding mistake ($72.3\%$).
 The first category includes issues like reentrancy, the second covers access control flaws, and the third involves coding mistakes such as arithmetic errors and logic absences.
 4.1.2. Benign Contracts
-To construct benign samples, we first collected a total of $796,437$ contracts deployed between June 2022 and June 2024 using Google’s BigQuery service.
-And we analyzed the transaction history of these contracts, focusing on the number of unique addresses interacted with the contracts from June 2022 to August 2024.
-The result showed that the adversarial contracts in our dataset often have very few unique users, with $355$ of them ($94.7\%$) interacting with only a single address, namely the attacker’s account, and a maximum of $6$.
+As contract-based DeFi attack detection approach is still a largely under-explored research direction, to date, there is no publicly available dataset that provides a comprehensive and representative set of ground truth benign contracts.
+Consequently, we need to build our benign dataset from scratch.
+Given the vast number and variety of smart contracts deployed on the blockchain, constructing a large-scale dataset with sufficient sample diversity and label accuracy requires extensive manual review efforts and is prohibitively time-consuming.
+In response to this, previous ML-based work Forta (For, 2024) follow a heuristic approach for collecting benign samples.
+Previous Approach.
+Forta assumes that all contracts verified on the chain explorer are benign.
+Despite being an intuitive design choice, analysis of our attack samples reveals that a small proportion of adversarial contracts are nonetheless verified.
+This may be because verification is an inexpensive task (often automated by development toolkits) and it has negligible impact on attack effectiveness, as existing detection methods do not focus on identifying adversarial contracts.
+Consequently, a more robust collection method is needed.
+Key Observations.
+A smart contract can be deployed either 1) through a normal transaction, where its to_address field is NULL with bytecode included in the input_data field, or 2) via an internal transaction (also known as a message call) initiated by an existing contract during its execution.
+Between June 2022 and June 2024, the number of contracts created via internal transactions is significantly higher than those created via normal transactions (approximately 20 times more).
+A prior research (Gupta and Shukla, 2019) that analyzed $1.16$ million contracts created via internal transactions found that these contracts were generated by only $9,228$ creator contracts, indicating that a small number of contracts were responsible for deploying a vast number of contract instances.
+Most of these instances originate from factory contracts, e.g., Uniswap’s factory contract  [^2] https://etherscan.io/address/0x5c69bee701ef814a2b6a3edd4b1652cb9cc5aa6f, which are designed to repeatedly create similar contracts.
+Additionally, we observe that all adversarial samples in our dataset are deployed through normal transactions.
+These contracts are typically invoked by a very limited number of unique addresses.
+Specifically, $355$ of them ($94.7\%$) interacting with only a single address, namely the attacker’s account, and a maximum of $6$ unique users.
 This low interaction count is consistent with the behavior of adversarial contracts, which typically feature access control and are designed for single-use only to attack specific victims.
-Based on this observation, we select contracts with at least $10$ unique interacting addresses as the determining factor for their benign nature.
-This filtering resulted in $210,643$ benign contracts.
+Our Method. Based on the observations, we construct our benign dataset following two key principles to ensure evaluation fairness, as well as dataset diversity and accuracy.
+  • To ensure fair evaluation, we focus solely on contracts created via normal transactions, consistent with the deployment patterns of all adversarial samples, while avoiding internal transactions that could drastically inflate the dataset with numerous near-identical instances.
+  • We select contracts with over a defined number of unique interacting addresses as an indicator of their benign nature, aiming to ensure reliability while maintaining sample diversity.
+We first query Google’s BigQuery service to collect contracts deployed via normal transactions between June 2022 and June 2024, identifying $796,437$ contracts.
+Then we analyzed their transaction history, focusing on the number of unique interacting addresses from June 2022 to August 2024, and retained contracts with at least $10$ unique interacting addresses, resulting in a final benign dataset of $210,643$ contracts.
+While no heuristic method can fully eliminate potential bias, our approach offers a pragmatic trade-off between dataset scale, diversity, and correctness.
 4.2. Feature Selection
-Based on empirical observations, we construct a selection of features for use in LookAhead’s ML model training.
-The set consists of features spanning over both the implementation and the deployment stages of a DeFi attack, aiming to capture any potential unexpected patterns of behaviour that belong to attackers.
-In this section, we explain our rationale behind the set construction.
+Based on empirical observations, we construct a selection of features for use in LookAhead’s ML model training and inference.
+The set consists of features spanning both the implementation and the deployment stages of a DeFi attack, aiming to capture any potential unexpected patterns of behaviour exhibited by attackers.
+In this section, we explain our rationale behind the choice of features and provide supporting statistical data whenever applicable.
 4.2.1. Deployment Features
@@
     Unknown: Fallback label for those that fail to fall into the above categories.
-Based on this taxonomy, we conduct a statistical analysis on the fund source of our contract samples.
-For our benign samples, more than three-quarter of them have Safe sources, approximately one-tenth of their labels are Anonymous.
-In contrast, the majority of the adversarial contracts have Anonymous fund sources, with only a small proportion having Safe labels.
+Statistical Data: In benign samples, we observe that more than $75\%$ of them have Safe sources, and approximately $10\%$ of them are funded through Anonymous sources.
+In contrast, more than $70\%$ of the adversarial contracts have Anonymous sources, with only about 10% having Safe labels.
 Transaction Data: $\{value,input\_data\_length,gas\_used\}$.
-Depending on the contract’s design, the contract creator has the ability to transfer native tokens to the contract upon deployment.
-The amount transferred is indicated by the $value$ attribute associated with the transaction.
+Depending on the contract’s design, the contract creator can transfer native tokens to the contract upon deployment.
 Adversaries typically do not perform such a transfer.
+To avoid the potential discrepancy across Ethereum and BSC, we binarize the value attribute to indicate whether such a transfer occurred (true or false), rather than using the exact amount.
 The input data field of a deployment transaction holds the contract’s bytecode, we measure its length to account for the size of the contract.
-Additionally, since the gas consumption of the deployment transaction is proportional to the complexity of the contract, we combine these metrics to effectively measure the scale of the contract.
+Additionally, we use $gas\_used$ to measure the computational resources consumed during deployment, which depends on the bytecode complexity and EVM execution, rather than gas fees, which relies not only on $gas\_used$ but also on the dynamically fluctuating gas price influenced by network congestion.
+Statistical Data: In our dataset, about 80% of benign contracts have input data exceeding 15,000 bytes, while over 50% of adversarial contracts fall below this threshold.
+To eliminate potential EVM-induced differences in $gas\_used$ between Ethereum and BSC, we simulate the deployment of all BSC adversarial samples on Ethereum using identical bytecode to obtain a consistent gas_used value.
+However, 24% of the contracts failed due to interactions with contracts absent on Ethereum.
+Among the successfully samples, 94% showed less than a 1% deviation compared to original value.
+Given this negligible difference, for the failed samples, we directly use their original gas_used value.
 Verification Status: $\{verified\}$.
@@
 Attackers, on the other hand, typically do not have the incentives to do so since it would expose their malicious intentions.
+Statistical Data: We observe in our dataset that only $6$ adversarial contracts ($1.6\%$) are verified , while more than $85\%$ of benign contracts are verified.
 4.2.2. Implementation Features
-Based on empirical observations, some of the characteristics that can be used to determine whether a contract is malicious or not include: implemented functions, internal and external function calls.
+Based on empirical observations, some of the characteristics that can be used to determine whether a contract is adversarial or not include: implemented functions, internal and external function calls.
 To capture other hidden or intrinsic properties of a contract, we also design a special technique called Pruned Semantic-Control Flow Tokenization (PSCFT) that provides a condensed representation of the behaviour of smart contracts.
@@
 Since adversarial contracts typically do not involve complex business logic, they often implement fewer functions.
-In our dataset, benign contracts frequently incorporate $\geq 20$ public functions, while adversarial ones tend to implement much fewer.
 We also calculate the number of flashloan callback functions and the proportion they take up over all public functions.
@@
 Consequently, they resort to utilizing flashloan to temporarily borrow a substantial amount of tokens and execute attack logic within corresponding callback functions.
-Since the size of lending pools are limited, attackers often borrow tokens from multiple sources, resulting in multiple flashloan callbacks.
-In our dataset, more than 60% of adversarial contracts implement flashloan callbacks, while almost all benign contracts do not involve such callbacks.
+Since the size of lending pools are limited, attackers may borrow tokens from multiple sources, resulting in multiple flashloan callbacks.
+Statistical Data: We can observe from the dataset that over 75% of benign contracts incorporate at least 20 public functions, whereas more than 75% of adversarial contracts implement 10 or fewer.
+Additionally, over 60% of adversarial contracts include flashloan callbacks, while more than 99% benign contracts involve no such callbacks.
 Function Calls: $\{token\_call\_count,max\_token\_call\_count,avg\_token\_call\_count,delegate\_c\\ all\_count,selfdestruct\_count\}$.
-We take into consideration external calls related to ERC-20 tokens, including their count and proportion.
+We take into consideration external calls related to tokens, including their count and proportion.
 Generally, the objective of DeFi attacks is to gain financial benefits, i.e., acquire valuable tokens.
 Therefore, external calls within adversarial contracts often involve token-related functions.
-The token-related functions we selected include, but are not limited to, common token standard (ERC, 2024a, b) functions and popular DEX (Uniswap, 2024; Adams et al., [n.
-d.]) functions.
-In our dataset, over $90\%$ of benign contracts do not have token-related calls.
-In contrast, almost all adversarial contracts involve token-related calls, with over two-thirds of them even having $10$ or more.
+The token-related functions we selected include, but are not limited to, common token standard (ERC, 2024a, b) functions and popular DEX (Adams et al., 2021) functions.
 Furthermore, we introduce two additional features to assess the behaviour of public functions in a smart contract.
 $max\_token\_call\_count$ is designed to capture the maximum number of token-related interactions within the execution flow of any public function of a contract, which helps us identify contracts with exceptionally high token interactions.
-Moreover, $avg\_token\_call\_count$ is also measured to demonstrate the average number of token-related calls in contracts.
+Moreover, $avg\_token\_call\_count$ is also measured to demonstrate the average number of token-related calls per public function.
 We also include the number of delegatecall instructions, as these are often implemented in proxy contracts.
 Additionally, some attackers have a tendency to self-destruct their contracts after completing the attack, so we also consider the presence of the selfdestruct instruction.
+Statistical Data: It can be observed from our dataset that over $90\%$ of benign contracts do not have token-related calls.
+In contrast, almost all adversarial contracts involve token-related calls, with over $70\%$ of them having $10$ or more.
+Figure 4. Side-by-side comparison of raw IR and PSCFT.
 Code Semantics: $\{PSCFT\}$.
 To understand the behaviours exhibited by DeFi attack adversaries and capture features that are potentially hidden from our empirical observations, we design a new representation for smart contracts called Pruned Semantic-Control Flow Tokenization ($PSCFT$).
-PSCFT improves upon the intermediate representation (IR) generated by Gigahorse to produce richer contextual information about the implementation of a contract.
-After constructing CFGs based on IR and going through an augmentation pipeline that retrieves and replaces function signatures and external addresses with appropriate semantic labels in the CFGs, PSCFT can effectively present a textual summary of function canonical names, control-flow graphs and external call chains.
-Figure 6.3 shows a contract PSCFT excerpt.
-One of the benefits of using PSCFT over IR is that IR alone does not provide useful information such as call target labels, and the size of the tokenization vocabulary is less manageable, rendering it not ideal for NLP training.
+In smart contracts, functionalities (including potential attack logic) are implemented through functions, and interactions with external entities, such as invoking the functions implemented in other deployed contracts, are performed via message calls.
+For adversarial contracts, such interactions (e.g., with victim contracts) are indispensable.
+Therefore, each function, along with its message calls and control flows, is capable of capturing the intrinsic behavioral characteristics of a contract, providing key information for determining whether the contract is adversarial or benign.
+To extract these critical features, we use Gigahorse (Grech et al., 2019) to lift the contract bytecode into an intermediate representation (IR), followed by a pruning process that remove extraneous elements while reconstructing control flows.
+To provide richer contextual information, the refined IR is fed into an augmentation pipeline that retrieves and replaces function signatures and external addresses with appropriate semantic labels.
+The resulting output is the fully PSCFT, which effectively represents a textual summary of function canonical names, control-flow graphs and external call chains.
+Figure 4 presents an excerpt comparing the raw IR of a contract with the corresponding PSCFT.
+We provide the detailed construction process of PSCFT in §5.2.
 5. Methodology
-Figure 6. LookAhead’s system building workflow
 In this section, we present our methodology for building the LookAhead system that performs adversarial contract classification. An overview of the complete framework is depicted in Figure 5.
 5.1. External Data Retrieval
-Some of the features used in our dataset are obtained from external sources, we identify related third-party services in the following section.
-Based the data, we develop algorithms such as fund source tracing to retrieve specific fields to be used for both training and inference of our ML models.
+Some of the features used in our dataset are obtained from external sources, we identify related third-party services in this section.
+Based the data, we use a algorithm to determine fund source for the contracts, enabling subsequent training and inference of our machine learning models.
 Transaction Data.
@@
 Prior to the practical deployment of LookAhead, we recommend setting up dedicated nodes for minimized data processing delays.
-We store and format contract deployment transactions and their associated attributes such as nonce, value, and gas consumption, along with the sender and input data properties for use in subsequent phases of our feature extraction pipeline.
+We store and format contract deployment transactions and their associated attributes such as nonce, value, and gas used, along with the sender and input data properties for use in subsequent phases of our feature extraction pipeline.
+Figure 5. LookAhead’s system building workflow
 Chain Explorer Data.
@@
 We describe the potential impact that this assumption could bring to the performance of LookAhead in §6.2.
-Furthermore, we develop a fund source labeling algorithm that recursively follow the transaction flow and trace all the way back to the original fund holder.
-Utilizing the address label cloud (lab, 2024) and public name tag provided by chain explorers, we manually review and assign fund source labels to each group of addresses.
-As an example, Tornado.Cash was given the label of Anonymous, while centralized exchanges like Coinbase were given the label of Safe.
+Furthermore, we develop a fund source labeling algorithm to determine a contract’s fund source.
+We first construct an address label dataset using the address label cloud (lab, 2024b, a) and the public name tags provided by chain explorers, following the definitions in §4.2.1.
+As an example, Tornado.Cash (Tor, 2024) is given the label of Anonymous.
+The algorithm traces fund flows from the contract deployer, recursively identifying the earliest funding source.
+If the source address has a known label, it is assigned; otherwise, tracing continues up to a predefined threshold depth.
+If no label is identified within this limit, the fund source of the contract is marked as Unknown.
 5.2. Pruned Semantic-Control Flow Tokenization (PSCFT)
-Based on state-of-the-art EVM contract binary lifter and analysis framework Gigahorse (Grech et al., 2019), we design a pipeline for 1) Constructing intra-contract Control Flow Graphs (CFG) and external call chains based on Intermediate Representation (IR), 2) Aggregating and extracting dataset features from EVM bytecode, and 3) Generating textual summaries for the smart contracts in our dataset.
+Based on EVM bytecode binary lifter framework Gigahorse (Grech et al., 2019), we design a pipeline for 1) Constructing intra-contract Control Flow Graphs (CFG) and external call chains based on Intermediate Representation (IR), 2) Aggregating and extracting dataset features from EVM bytecode, and 3) Generating textual summaries for the smart contracts in our dataset.
 5.2.1. Implementation Feature Extraction
-To enable the extraction of implementation features (cf.
-§4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse.
-First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and $selfdestruct\_count$.
-Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers), with the help of PublicFunctionSelector, which maps public functions to their selectors, to determine the total number of public and private functions.
+To enable the extraction of implementation features (see §4.2.2) simultaneously with the generation of the IR, we implemented a custom-designed Souffle Datalog program based on existing rules from Gigahorse.
+1) First, by utilizing the Statement_Opcode rule, which maps each statement to its corresponding opcode, we can directly extract the number of occurrences of critical opcodes, such as DELEGATECALL and SELFDESTRUCT, which corresponds to features $delegate\_call\_count$ and selfdestruct_count.
+2) Next, we leverage the FunctionEntry rule, which contains all function entry points (identifiers), with the help of PublicFunctionSelector, which maps public functions to their selectors, to determine the total number of public and private functions.
 This combination also helps identify flashloan callback functions.
-Finally, to analyze the external function calls within a contract, we design a rule called ExternalCallResolved.
-This rule provides details about all resolved external calls made within the contract, including their selectors and signatures.
-Additionally, we implemented logic to handle special cases, such as identifying fallback functions and transfers of native tokens, based on the arguments and values passed to the calls.
+3) Finally, to analyze the external function calls, we design the ExternalCallResolved rule, providing details about all resolved external calls made within the contract, including their selectors and signatures.
 With the help of ExternalCallResolved, we can keep track of token-related calls and their related information including their total count, maximum occurrences, and average usage.
 5.2.2. CFG Construction and Pruning
-Based on the IR of a contract, we construct a function-by-function visualization $F=\{F^{private}\cup F^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements.
+Based on the IR of a contract, we construct a function-by-function visualization =$\mathcal{F}=\{\mathcal{F}^{private}\cup\mathcal{F}^{public}\}$ of the internal structure of contract methods, each consisted of control flow blocks and programming statements.
 Functions have 4-byte selector derived from the Keccak hash (SHA3) of their canonical names and can be either private or public.
-Each function contains multiple basic blocks (BB) chained together ($\{\mathcal{F}_{i}=\{BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|\mathcal{F}_{i}|}\}\}$), and each block contains multiple statements.
+Each function contains multiple basic blocks (BB) chained together ($\{F_{i}=(BB^{i}_{0},BB^{i}_{1},\cdots,BB^{i}_{|F_{i}|})\}$), and each block contains multiple statements.
 Blocks and statements are identifiable via unique numerical IDs, for example, $90x11b$ 9.
-The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B=\{\mathcal{BB}^{i}_{j}:\mathcal{BB}^{i}_{j}\in\mathcal{F}_{i}\}$ where $\mathcal{F}_{i}\in F$, and $(\mathcal{BB}^{i}_{x},\mathcal{BB}^{i}_{y})$ is an arc in $A_{i}$ if $\mathcal{BB}^{i}_{x}$ is a predecessor of $\mathcal{BB}^{i}_{y}$ and $\mathcal{BB}^{i}_{y}$ is a successor of $\mathcal{BB}^{i}_{x}$.
+The relationship between two blocks can be discovered by following the predecessor and successor IDs associated with each block, each block can have multiple predecessors and successors, hence we can create a directed graph $G_{i}=(B_{i},A_{i})$ such that the set of nodes $B_{i}=\{BB^{i}_{j}:BB^{i}_{j}\in F_{i}\}$ where $F_{i}\in\mathcal{F}$, and $(BB^{i}_{x},BB^{i}_{y})$ is an arc in $A_{i}$ if $BB^{i}_{x}$ is a predecessor of $BB^{i}_{y}$ and $BB^{i}_{y}$ is a successor of $BB^{i}_{x}$.
 Statements could be constant/variable declarations, internal/external function calls, function terminations, memory copies, storage initialization, etc., we focus on constructing a CFG based on function call statements.
@@
 Gigahorse can only recover the names for some of the public functions and external call methods, hence if the derivation of canonical names from function selectors failed, we will leave them as-is and recover their names in the semantic recovery stage.
-For any private function $\mathcal{F}^{private}_{j}\in F^{private}$, we give it a canonical name of $InternalFunction_{i}$.
-Next, we remove all statements not related to function calls from every block $\mathcal{BB}^{i}_{j}\in B_{i}$.
-Then, we perform a depth-first search (DFS) over $G$ to remove any block $\mathcal{BB}^{i}_{k}$ that contains no statements (i.e.
-$|\mathcal{BB}^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours.
+For any private function $F^{private}_{j}\in\mathcal{F}^{private}$, we give it a canonical name of $InternalFunction_{i}$.
+Next, we remove all statements not related to function calls from every block $BB^{i}_{j}\in B_{i}$.
+Then, we perform a depth-first search (DFS) over $G$ to remove any block $BB^{i}_{k}$ that contains no statements (i.e.
+$|BB^{i}_{k}|=0$) from $B_{i}$, along with its incoming and outgoing arcs from $A_{i}$, and update the the predecessor and successor IDs for its neighbours.
 Finally, we assign canonical names $BB^{i}_{j},j\in[0,|B|-1]$ to basic blocks that are still left in $B_{i}$.
 5.2.3. Semantic Recovery
-For an external call statement like the one shown in section PSCFT and Feature Construction in Figure 6, we aim to recover: 1) An address label for 0xBA12222222228d8Ba445958a
-75a0704d566BF2C8
-, 2) A human-readable name for the target method 0x5c38449e.
-Similar to our fund source labelling algorithm, we attempt to match the target address with an address associated with a known label in Etherscan’s address label cloud and public name tags.
+For an external call statement like the one shown in §PSCFT and Feature Construction in Figure 5, we aim to recover: 1) An address label for 0x5777d92f208679db4b9778590
+fa3cab3ac9e2168
+, 2) A human-readable name for the target method 0x490e6cbc.
+Similar to our fund source labelling algorithm, we consider to match the target address with an address associated with a known label in chain explorers’ address label cloud and public name tags.
 For contracts that cannot be matched with an existing label, we check if they are open source, when they are, we use their contract name as the label.
@@
 However, thanks to the deterministic nature of hash functions, third-party service such 4byte (4by, 2024) are able to construct and maintain a large database of publicly known hash-to-name mappings.
-Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we extend Gigahorse’s name recovery ability with 4byte’s service by sending requests to their APIs upon failed recovery by Gigahorse.
+Therefore, in hopes to include as much function names and as little unknown functions in the PSCFT as possible, we attempt to extend Gigahorse’s name recovery ability with 4byte’s service by sending a request to their API upon failed recovery by Gigahorse.
 If the recovery still fails, we fallback to the default label UnknownFunc.
-If the aforementioned process was completed successfully, the external call statement will be converted to Balancer.flashLoan(…args).
-And we can produce a textual summary for smart contracts containing control-flow and semantic call information simply by concatenating $\mathcal{F}_{i}=\{\mathcal{BB}^{i}_{0},\cdots,\mathcal{BB}^{i}_{|\mathcal{F}_{i}-1|}\}\in F$ together in the form shown in Figure 6.3, Semantic Recovery.
+If the aforementioned process was completed successfully, the external call statement will be converted to UniswapV3.flash(…args).
+And we can produce a textual summary for smart contracts containing control-flow and semantic call information simply by concatenating $F_{i}=(BB^{i}_{0},\cdots,BB^{i}_{|F_{i}-1|})$ $\in\mathcal{F}$ together in the form shown in Figure 4.
 5.3. Classifier Design
-5.3.1. Semantic Transformer
+5.3.1. Transformer
 We train a transformer-based Natural Language Processing (NLP) model $\mathcal{M}^{transformer}$ solely on the PSCFT constructed from the contracts in our dataset to learn the semantics and behaviours of the implementation of both benign and adversarial smart contracts.
@@
 To ensure the model’s generalization ability and improve training efficiency, we implement regularization techniques such as early-stopping and dropouts to prevent model over-fitting.
-Overall, this transformer model provides a semantic view of DeFi smart contracts.
+Overall, this transformer model provides a semantic view of smart contracts.
 5.3.2. Ensemble Classifier
-Aiming to embed the features selected based on our empirical observations into the semantic view of the transformer model, we design an ensemble classifier that merges the classifier models trained on basic features with the transformer model trained on PSCFT together to produce a meta classifier model with better predictive performance.
+Aiming to embed the features selected based on our empirical observations into the semantic view of the transformer model, we design an ensemble classifier that merges the classifier models trained on empirical features with the transformer model trained on PSCFT together to produce a meta classifier model with better predictive performance.
 Candidate Classifiers.
 Once all features have been extracted, we can proceed to train a supervised learning classifier using a variety of ML algorithms, outputting multiple candidate models for further evaluation.
-These models include classical approaches such as Logistic Regression $\mathcal{M}^{classifier}_{LR}$ and Decision Trees $\mathcal{M}^{classifier}_{DT}$, ensemble learning methods such as bagging-based Random Forest $\mathcal{M}^{classifier}_{RF}$ and XGBoost $\mathcal{M}^{classifier}_{XGBM}$ based on boosting.
+These models =$\mathcal{M}=\{M^{classifier}_{LR},M^{classifier}_{DT},M^{classifier}_{RF},M^{classifier}_{XGBM}\}$ include classical methods such as Logistic Regression and Decision Trees, ensemble learning methods such as bagging-based Random Forest and XGBoost which is based on boosting.
 Due to the limited quantity of adversarial samples, to mitigate the bias introduced by class imbalance, we employ the ADASYN oversampling method (He et al., 2008) that synthesizes additional samples to bring the number of adversarial samples closer to that of the benign ones in the training dataset.
-For the fund source feature, which consists of categorical data (ref.
-§4.2.1) with low cardinality, we apply one-hot encoding.
+For the fund source feature, which consists of categorical data (see §4.2.1) with low cardinality, we apply one-hot encoding.
 To facilitate better comparability among input values in numerical terms of varying dimensions, we standardize the dataset using Z-score normalization.
 This process brings each feature value around to the centre of 00 with a standard deviation of $1$.
-After training has completed for each of the candidate classifiers, we evaluation their performance using metrics such as Recall and F1-Score to determine an optimal model $\mathcal{M}^{classifier}_{optimal}\in\{\mathcal{M}^{classifier}_{LR},\cdots,\mathcal{M}^{classifier}_{XGBM}\}$ for the meta classifier training and inference in the next stage.
+After training has completed for each of the candidate classifiers, we evaluate their performance using a function $f:\mathcal{M}\rightarrow\mathbb{R}$ that incorporates metrics such as Recall and F1-Score to determine an optimal model $M^{classifier}$ such that $M^{classifier}=\text{argmax}_{M\in\mathcal{M}}f(M)$ for the meta classifier training and inference in the next stage.
 Meta Classifier.
-Using the stacked generalization (stacking) machine learning algorithm, we design a blender that stacks $\mathcal{M}^{transformer}$ and $\mathcal{M}^{classifier}_{optimal}$ together to form a higher level model $\mathcal{M}$ known as a meta classifier.
+Using the stacked generalization (stacking) machine learning algorithm, we design a blender that stacks $M^{transformer}$ and $M^{classifier}$ together to form a higher level model $M^{meta}$ known as a meta classifier.
 By feeding the predictions produced by the two base models into the meta model, we are able to 1) Combine the strengths of each model and automatically learn the optimal way to construct the inference result based on model significance, 2) Optimize the final classification boundaries to achieve more accurate detection of adversarial contracts.
-We use K-Nearest Neighbors (KNN) to as our meta classifier $\mathcal{M}$ due to its simplicity, effectiveness, and ability to make predictions based on the proximity of data points in the feature space.
-KNN is well-suited for combining the outputs of different classifiers such as the transformer model and one of the candidate classifiers, because it is able to combine the complementary strengths of both models by analyzing their predictions and finding the optimal decision boundary.
+Given that the meta classifier operates on low-dimensional inputs (only the two probability values between 00-$1$ outputted from the $M^{classifier}$ and $M^{transformer}$, indicating the likelihood of a contract being adversarial) and without incorporating any additional information about the contracts, employing a simple model is both sufficient and effective.
+This not only ensures efficiency but also mitigates the risk of overfitting by limiting model complexity.
+Accordingly, we only consider simple models, including K-Nearest Neighbors (KNN), Logistic Regression (LR), Support Vector Machines (SVM), and Decision Trees (DT), and evaluate their performance to select the most suitable meta classifier in experiments.
 6. Evaluation
 We implement LookAhead in Python using open source machine learning frameworks PyTorch and Tensorflow to achieve both CPU and GPU-accelerated training and inference of our classifier models.
-All experiments are conducted on a server running a 3.70GHz Intel Xeon processor with 32 threads and a NVIDIA A100 GPU with 40GB of VRAM.
+All experiments are conducted on a server running a 3.70GHz Intel(R) Xeon(R) processor with 32 threads and a NVIDIA A100 GPU with 40GB of VRAM.
 6.1. Experiment Setup
-6.1.1. Dataset
+6.1.1. Datasets
 Token and proxy contracts are two of the primary types of contracts used on blockchains, while they are commonly used by developers, most adversaries do not rely on those types of contracts for their core attack logic.
@@
 By analyzing whether the contracts match common standards, we found that in the dataset we built, $184,386$ benign samples were token contracts, $10,037$ were proxy contracts, and none of the adversary contracts belonged to these two categories.
-Therefore, we remove them from both our training and testing dataset the same way as it was in previous system Forta (For, 2024).
+Therefore, we remove them from both our training and testing dataset the same way as it was in previous work Forta (For, 2024).
 A distinguishing aspect of our research when compared to traditional supervised machine learning problems is the explicit chronological order of contract deployment.
@@
 One of LookAhead’s objectives is to provide a ML classifier model that is pre-trained on past DeFi incidents but capable of detecting potential future attacks.
-Consequently, traditional cross-validation techniques, which assume input values are independent with respect to each other, are not suitable for our evaluation.
-Instead, we use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing.
+Consequently, we sort the contract samples chronologically and use the earliest $80\%$ of the contracts for training and reserve the latest $20\%$ for testing.
 Out of the $80\%$ of contracts used as training data, we take the last $25\%$ of most recent samples as the training set for the meta classifier, and the rest for the transformer and candidate models.
 6.1.2. Research Questions
+Table 1. Performance breakdown of different models.
 Model Accuracy Precision Recall F1-Score FPR
@@
 XGBoost 0.9932 0.8333 0.8784 0.8553 0.0041
-NLP Classifier Transformer 0.9876 0.7429 0.7027 0.7222 0.0057
-Forta - 0.8778 0.5520 0.6208 -
-Meta Classifier KNN 0.9953 0.9286 0.8667 0.8966 0.0016
-Table 1. Performance breakdown of different models.
+NLP Classifier Forta - 0.8778 0.5520 0.6208 -
+Transformer 0.9876 0.7429 0.7027 0.7222 0.0057
+Meta Classifier LR 0.9950 0.9683 0.8133 0.8841 0.0006
+DT 0.9913 0.7901 0.8533 0.8205 0.0054
+SVM 0.9947 0.9143 0.8533 0.8828 0.0019
+KNN 0.9953 0.9286 0.8667 0.8966 0.0016
 We attempt to address the following research questions:
   1. 1)
-    RQ1. How effective is LookAhead’s system of classifiers in detecting the adversarial contracts used in DeFi attacks, especially when compared to previous state-of-the-art solutions?
+    RQ1. How effective is LookAhead in detecting the adversarial contracts used in DeFi attacks?
   2. 2)
-    RQ2. How efficient can the LookAhead system be in identifying adversarial contracts, can it produce a valid response quick enough so that defense actions could take place?
+    RQ2. How do different empirical features contribute to model decision-making and how does PSCFT improve performance compared to using raw IR?
   3. 3)
-    RQ3. How practical is it for individuals or organizations to deploy the LookAhead system and what does the performance of LookAhead look like in real-world scenarios?
+    RQ3. How efficient can LookAhead be in identifying adversarial contracts?
+  4. 4)
+    RQ4. How practical is LookAhead  in real-world scenarios?
 6.2. RQ1: Effectiveness
@@
 In addition to commonly used binary classification performance metrics such as Accuracy, Precision, Recall and F1-Score, we also calculate False Positive Rates (FPR), which holds particular significance in our scenario.
-6.2.2. Experiment Results
-Table 1 presents our experiment results for evaluating the performance of LookAhead’s ML models.
-We calculate the performance metrics of various models used in the process of forming the LookAhead detection system, including $4$ candidate classifiers, a transformer-based classifier, and a meta classifier.
-Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest (RF) has the lowest recall ($0.7432$).
-Overall, the XGBoost-based classifier outperforms others, demonstrating the best performance across most metrics compared to other candidates.
-Its false positive rate reaches as low as $0.0041$, and it exhibits a high effective detection rate for adversarial samples.
+6.2.2. Evaluation Results
+Table 1 presents our experimental results evaluating the performance of various ML models (§5.3) for the LookAhead system, including four candidate classifiers, a transformer-based classifier, and four meta classifiers.
+Among the candidate classifiers, XGBoost achieves the highest F1-score ($0.8553$), while Random Forest has the lowest recall ($0.7432$).
+Overall, the XGBoost-based classifier outperforms other candidate classifiers, demonstrating the best performance across most metrics compared to others.
+It achieves a false positive rate as low as $0.0041$ while maintaining a high effective detection rate for adversarial samples.
 Other models also perform well.
-Except the one based on Decision Tree (DT), the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method.
-The meta classifier, K-nearest Neighbors (KNN), combines the optimal candidate model XGBoost with the transformer to achieve the best overall performance with the highest accuracy ($0.9953$), F1-Score ($0.8966$), and the lowest FPR ($0.0016$).
-6.2.3. Comparison with the Baseline
-There have been studies that analyze smart contracts to detect contract vulnerabilities, but to the best of our knowledge, this is the first academic work that focus on analyzing smart contracts for the detection of adversarial contracts used by attackers across all types of DeFi attacks.
-The goal of both research directions is to prevent attackers from exploiting vulnerable contracts for profits, the only different is that our method aim to stop attacks from causing damage to potential targets while other works help potential targets reduce the possibility of being attacked.
-However, we emphasize that existing works on detecting vulnerabilities often focus on specific types of attacks/vulnerabilities, offering limited protection capabilities.
-Therefore, when comparing our results with the baseline, our goal is to: 1) Highlight the general applicability of LookAhead’s system in detecting all types of DeFi attacks, 2) Demonstrate the performance of LookAhead against previous state-of-art solutions.
+Except the one based on DT, the F1-scores of all other classifiers exceed $0.79$, proving the validity of our method.
+Building upon the results of the candidate models, different meta classifiers were constructed by combining the optimal candidate model XGBoost with the transformer.
+The results show that three of the four meta classifiers achieved higher F1-scores compared to the standalone XGBoost classifier.
+Among them, the KNN-based classifier demonstrated the best performance, achieving the highest F1-score ($0.8966$) while maintaining a low FPR ($0.0016$).
+6.2.3. Comparison with Previous Works
+We compare our approach with two state-of-the-art baselines: Forta (For, 2024), a NLP-based method for detecting all adversarial contracts, and BlockWatchDog (Yang et al., 2024), a specialized work focused exclusively on detecting reentrancy attack contracts.
 Universality.
-We compare the performance of LookAhead with a previous state-of-the-art academic work (Yang et al., 2024) that specifically detects reentrancy attack contracts to demonstrate the universality of the LookAhead system.
-When our trained meta classifier model was tested on all the reentrancy related adversarial contracts in our test dataset, we observed that LookAhead raised an alert for $83.33\%$ of the attack contracts, which is very similar to the results presented by (Yang et al., 2024).
-This observation along with the experiment results presented in §6.2.2 show that despite being trained on contracts used in a variety of attacks of different types (cf.
-§4.1.1), LookAhead still achieves comparable performance in detecting any specific types of DeFi attacks.
+To evaluate the universality of the LookAhead system, we compare its performance against BlockWatchDog, which specifically targets reentrancy attack contracts.
+We test our meta classifier on all the reentrancy related adversarial contracts in our test dataset, we observe that LookAhead successfully identified $83.33\%$ of them, which is on par with the results presented by (Yang et al., 2024).
+This observation along with the experiment results presented in §6.2.2 show that LookAhead not only generalizes well across diverse types of DeFi attacks but also maintains competitive performance when applied to specific attack categories.
 Performance.
-The only similar work that uses machine learning for adversarial contract detection that was made public so far is the Forta Network (For, 2024).
-However, their approach relies solely on EVM bytecode while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts.
+The only similar work that uses machine learning for adversarial contract detection publicly available so far is the Forta Network (For, 2024).
 We use Forta as the baseline for comparing the performance of our ML models.
-We consider Forta an NLP-based method, hence we perform a direct comparison between their model and our transformer model first, thereafter we measure how much benefits can our meta classifier approach provide over Forta’s implementation.
-As the results in Table 1 indicate, our transformer alone can perform better and achieve higher F1-Socre than Forta, demonstrating the advantages of our Pruned Semantic-Control Flow Tokenization design.
-Our meta classifier pushes the boundary even further by combining the capabilities of transformer model with the capabilities of other candidate classifiers to achieve F1-Score as high as $0.8966$.
-Furthermore, as described in §5.1, contract verification status feature introduces some bias into our dataset, to provide a fair comparison, we also evaluate the performance of our models with $verified$ feature removed.
-Experiments show that even with the feature $verified$ removed from our training and evaluation dataset, our classifiers still significantly outperform the most recent model published by Forta (released on February 6, 2023), achieving an F1-Socre of 0.8696, highlighting the superiority of LookAhead’s solution.
-6.3. RQ2: Efficiency
-Efficiency testing evaluates LookAhead’s resource consumption during its operating time, such as processing power requirements, memory usage, and detection throughput.
-All these contribute to delays in the identification of attackers and potential victims.
-Given the frequency of smart contract deployment transactions in decentralized finance, LookAhead should be able to quickly respond and react to potential attack activities within typical rescue time frames.
-The most time-consuming step in LookAhead’s detection pipeline is the the contruction of PSCFT.
-Thanks to the rapid advancement of ML, classifier inference can be done on either CPUs or GPUs in a batched manner, which means in most cases it is a significantly faster process than binary lifting and analysis.
-Additionally, for external data augmentation pipelines that require sending API requests to outside sources, we can often easily employ multi-threading techniques to achieve high data retrieval concurrency.
-Thus, it can be assumed that the time required for analyzing a smart contract is roughly equivalent to the time needed for constructing its PSCFT.
-To evaluate the efficiency, we measured the analysis time for all adversary contracts.
-The results showed that, except for one large contract with bytecode close to the maximum size, which took about $455$ seconds, most other contracts’ PSCFTs were constructed in under $20$ seconds, with an average time of $7.34$ seconds.
-Additionally, by considering the rescue time frames (cf.
-§3.2.4) of their corresponding DeFi attacks, we found that in $92.8\%$ of the cases, our system could respond in time by detecting the adversary contract before the adversary transaction occurred, leaving plenty of room for defense actions to be taken by affected parties and proving the viability of the LookAhead system.
-6.4. RQ3: Practicality
-We consider the practicality of the LookAhead system crucial to the validity of our research, therefore we conducted experiments measuring the detection performance of LookAhead to verify its effectiveness in realistic, dynamic environments.
-From August 14 to August 18, 2024, we deployed the LookAhead system to a cloud server and conducted a live experiment to measure LookAhead’s performance in real-world scenarios.
-We use a meticulous review process that require us to 1) Wait for $2$ weeks (September 1) to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
-Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, in which we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks.
-These attacks caused over $50,000$ USD in financial losses.
-In terms of detection efficiency, except only one case where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction.
-Additionally, after thorough analysis, we observed that almost all the false positives are related to arbitrage, showing that the behaviour exhibited by arbitrage trading resembles some patterns in adversarial activities.
-For example, the dependency on flashloans, the high number of token ralated function calls, etc.
+Forta is an NLP-based method that relies solely on EVM bytecode, while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts.
+We perform a direct comparison between their model and our transformer model first, then we measure how much benefits can our meta classifier approach provide over Forta’s implementation.
+As the results in Table 1 indicate, our transformer alone can achieve much higher F1-Score than Forta, demonstrating the advantage of our PSCFT design.
+LookAhead’s meta classifier pushes the boundary even further by combining the capabilities of transformer model with the capabilities of other candidate classifiers to achieve an F1-Score of $0.8966$.
+As described in §5.1, contract verification status feature introduces some bias into our dataset, to provide a fair comparison, we also evaluate our models with $verified$ feature removed.
+The results show that even without this feature, our classifiers still significantly outperform the most recent model published by Forta (released on February 6, 2023), achieving an F1-Socre of $0.8696$, highlighting the effectiveness of our solution.
+6.2.4. Time Series Cross-Validation
+As outlined in 6.1.1, our evaluation respects the chronological order of contract deployment, accounting for both the evolving nature of adversarial behaviours and LookAhead’s design to detect future attacks based on historical data.
+Table 1 shows the model performance results obtained by using the entire dataset with a 4:1 chronological split.
+We identify XGBoost as the optimal candidate classifier and KNN as the best-performing meta model.
+To validate the most appropriate candidate and meta models for our system and ensure its robustness under varying data conditions, we conduct evaluations using expanding window cross-validation.
+During this process, the dataset is chronologically divided into five equal splits.
+The training window starts with the first split and progressively expands by including additional splits, while the subsequent split serves as the test set in each iteration.
+After cross-validation, we conclude that XGBoost achieves the highest average F1-score (0.8173), demonstrating stable performance across varying data splits.
+While it did not always outperform others in every split, its overall consistency reaffirms its suitability as the optimal candidate classifier for our system.
+For the meta classifiers, which combine the optimal candidate model XGBoost with the transformer, the KNN-based model achieves the highest average F1-score ($0.8355$).
+However, its advantage over other models were marginal, suggesting that the choice of meta classifier models has limited impact on overall performance.
+6.3. RQ2: Interpretability
+6.3.1. Empirical Features
+Understanding the weight distribution of features is useful for gaining insights into the underlying decision-making process of the models.
+To this end, we apply the interpretability method SHAP (Shapley Additive exPlanations) (Lundberg and Lee, 2017) to analyze the importance of empirical features for each candidate model.
+The results reveal that both deployment and implementation features play significant roles in guiding the models to a decision.
+In the best-performing XGBoost model, the top three most important features are avg_token_call_count, verified and fund_from_Anonymous, all of which also ranked within the top 10 across other candidate models.
+Additionally, balanceOf_call_count, fund_from_Safe and func_count consistently appear within the top 15 in at least three candidate models, indicating their stable contribution to the model’s decision-making process.
+The notable importance of these features may be attributed to their strong correlation with typical attack patterns, allowing the model to effectively capture behavioural characteristics of adversarial contracts.
+6.3.2. PSCFT
+The design of PSCFT abstracts the IR of a contract into a more concise and semantically meaningful format.
+To concretely demonstrate how this design enhances model performance, we compare the transformer classifier trained on PSCFT with one trained on the raw IR generated by Gigahorse.
+The results show a significant performance gap, where the model using raw IR achieves an F1-score of only 0.4068, substantially lower than the 0.7222 achieved by the PSCFT-based model.
+This performance discrepancy highlights the limitations of directly using raw IR, which contains numerous low-level instructions (e.g., arithmetic operations, memory reads/writes) and excessive variables that obscure critical call patterns and their control flows.
+Such extraneous details increase input complexity, making it harder for the model to identify meaningful behavioural patterns relevant to adversarial contract.
+Our pruning process can effectively remove redundant instructions and enable semantic recovery, allowing PSCFT to focus on critical external calls and control flows, leading to clearer and more reliable model decisions and improved prediction accuracy.
+6.4. RQ3: Efficiency
+LookAhead  aims to detect adversarial contracts and enable rapid responses within the rescue time frame (see §3.2.4). To this end, we evaluate the efficiency using all adversarial samples.
+We employed a linear detection pipeline, which processes each sample through the following steps: 1) retrieve deployment features via external APIs, 2) execute our custom-designed Souffle Datalog and PSCFT generation scripts to extract implementation features, and 3) preprocess the data and feed it into pre-trained XGBoost and transformer classifiers.
+Their outputs are then combined by a pre-trained meta-classifier for the final prediction.
+We measure the time taken at each step.
+The results showed that, on average, external API requests took $3.78$ seconds, while data preprocessing and model inference required only $0.01$ seconds.
+The most time-consuming step is binary lifting and analysis.
+Except for one large contract (taking $458$ seconds), most PSCFTs were generated under $20$ seconds, averaging $7.52$ seconds.
+Furthermore, by considering the rescue time frames of the corresponding DeFi attacks, we found that in $92.8\%$ of the cases, our system successfully detected the adversarial contract before the adversarial transaction occurred.
+Notably, efficiency can be further improved by employing multi-threading and local blockchain nodes, which could substantially reduce the time required to obtain deployment features.
+6.5. RQ4: Practicality
+Setup We conducted live experiments targeting the Ethereum mainnet across two time periods, 14–18 August 2024 and 23 January–9 February 2025, by deploying the LookAhead system to the server.
+The system continuously monitored the transactions in each new block via RPC providers (Alc, 2024) in near real-time.
+Upon detecting a contract creation (through a normal transaction), scripts were executed to extract features, which were then preprocessed and fed into the pre-trained classifiers (XGBoost and KNN, the best-performing setup as identified in RQ1).
+If classified as adversarial, the contract address was recorded for further verification.
+Review Process We use a meticulous review process as follows to verify the outputs:
+  1. 1)
+    Wait for $2$ weeks after each experimental period to ensure attacks have been confirmed;
+  2. 2)
+    Manually inspect follow-up transactions involving the contracts to validate adversarial nature;
+  3. 3)
+    Analyze and record key data, e.g. the financial damage caused by confirmed attacks, the time window between contract deployment and corresponding adversarial transaction.
+Table 2.
+Representative true positive adversarial contracts detected during the live experiment, involved in attacks triggered by different root causes.
+Window indicates the time interval between contract deployment and the adversarial transaction.
+Runtime denotes the detection time, and Victim specifies whether the victim address was exposed during contract creation.
+Address Time Window Runtime Victim Root Cause Loss (\$)
+0x9f2790 2024/8/16 25344s 11.53s No Price Manipulation 15.5k
+0x90744c 2024/8/16 636s 4.84s Yes Stale Oracle 21.5k
+0x152619 2024/8/18 336s 5.67s Yes Access Control Flaw 1.2k
+0x172133 2025/1/23 3504s 6.12s No Absence of Input Validation 2.5k
+0x203f20 2025/2/6 120s 14.35s Yes Arbitrary Transfer 37k
+0xbfe102 2025/2/6 96s 6.82s Yes Reentrancy 8.7k
+Results During the experiment period, LookAhead analyzed a total of $23,463$ contracts and flagged $129$ as adversarial.
+Among these, $62$ involved follow-up transactions, in which we finally found $26$ true positives that were used to perform attacks, resulting in over \$150,000 USD in financial losses.
+These attacks were driven by various underlying issues, with Arbitrary Transfer (Ma et al., 2023) being the most common, involving 6 contracts.
+Table 2 presents representative adversarial contracts involved in attacks stemming from different root causes, highlighting the universality of our approach, which effectively identifies various types of adversarial behaviours.
+In terms of detection efficiency, except only two cases where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction.
+Additionally, we observed most of the false positives are related to arbitrage, showing that arbitrage behaviour resembles some patterns in adversarial activities, such as reliance on flashloans and frequent token-related function calls.
+To evaluate potential false negatives, we reviewed the alerts posted by security company (e.g.
+(blo, 2024; cer, 2024; slo, 2024)) during the experiment period, as manually checking for false negatives among a large number of contracts is challenging.
+This process identified four reported attacks that fell within our detection scope (§3.2.1), all of which had already been correctly flagged.
+The fact that, during our live experiment, LookAhead was able to respond in a timely manner before the majority of real-world attacks occurred, and successfully detected all externally reported attacks, further demonstrates its practicality in real-world scenarios.
 7. Discussion
-Threats to Validity.
-Limited by the API responses given by the blockchain explorers, we can only go as far as pinpointing the date of contract verification without knowing the exact time.
-Thus we cannot determine whether verification was done at the time of deployment or not.
-Additionally, due to the scarcity of actual adversarial contract samples, the ratio between the two types of samples in our final dataset is approximately $1:40$.
-This imbalance could affect the performance of our models.
-Furthermore, benign contracts were selected based on empirical observations and the assumption that adversarial contracts typically have fewer unique interactions with other accounts.
-There is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results.
-Smart contracts can be created through internal transactions triggered during the execution of a normal transaction.
-Those contracts are not included in our dataset.
-Limitations.
-Our current method has certain limitations in terms of the spectrum of adversarial contracts that LookAhead is capable of detecting.
-When selecting features, we rely on existing attack patterns seen in DeFi attacks involving financial losses and those in prevalent attack scenarios.
-Consequently, there is a probability that we might have missed attacks that do not result in financial damages and certain adversarial contracts with novel attack vectors.
-Furthermore, in the cat-and-mouse game of DeFi attack and defense, attackers may evade our detection by using techniques like code obfuscation to alter the CFG of contract programs or by utilizing KYC-required exchanges as funding sources that makes it harder for our system to recognize adversarial intentions.
+7.1. Threats to Validity
+1) Limited by the APIs given by the blockchain explorers, we cannot determine whether contract verification occurred upon deployment; 2) Our dataset suffers from a class imbalance of approximately $1:40$ due to the scarcity of adversarial contracts, which may affect model performance even after oversampling; 3) Benign contracts were selected based on heuristic approach.
+There is still a possibility that we have mislabelled some adversarial contracts as benign ones; 4) As discussed in §4.1.2, our study does not consider contracts deployed via internal transactions.
+The inherent differences between the two deployment methods may affect certain deployment features, introducing potential bias to our classifier when detecting contracts deployed via internal transactions.
+7.2. Limitations
+In addition to the inherent limitations of our threat model definition (§3.2), there are mainly two aspects of LookAhead’s design that could potentially hinder its performance.
+Detection Evasion Since our classifier relies on extracted features and historical attack patterns, it remains susceptible to evasion techniques.
+In the cat-and-mouse game of DeFi attack and defense, attackers may adapt their strategies to bypass our feature-based and model-driven approach.
+For instance, they could adopt code obfuscation to alter the contract’s CFG or utilize KYC-required exchanges as funding sources, making it harder for our system to recognize adversarial intentions.
+NLP-based Approach LookAhead’s NLP-based models may not make full use of the control flow graph structure present in smart contract code.
+More advanced ML architectures such as Graph Neural Network (GNN) could be used in future works in combination with a more intricate representation design that capture both intra-function CFGs and inter-function call relationships.
 8. Related Work
-There have been many studies concerning the detection of abnormal contracts, primarily targeting Ponzi schemes and honeypots.
-They share similarities in methods which could be categorized into code similarity detection (Bartoletti et al., 2020; Sun et al., 2020), machine learning (Chen et al., 2020; Hu et al., 2022; Ji et al., 2023; Zheng et al., 2023), and symbolic execution (Chen et al., 2021).
-However, to the best of our knowledge, this is the first academic work that focuses on identifying adversarial contracts, and the first to support detection across all types of DeFi attacks.
-Forta (For, 2024) also proposed a method that uses machine learning techniques to identify adversarial contracts, but their approach is limited to bytecode-level analysis.
-Aside from Forta, the only similar work is BlockWatchDog proposed by Yang et al.
-(Yang et al., 2024), which uses cross-contract static dataflow analysis technique to detect adversarial contracts in order to identify reentrancy vulnerabilities.
-However, their focus is solely on reentrancy attacks, which account for only $6.7\%$ of the attacks in our dataset, and their method relies heavily on knowing the victim addresses in advance.
+Adversarial Transaction Detection.
+Prior research primarily focus on detecting adversarial transactions used in DeFi attacks.
+Some works (Chen et al., 2020; Grossman et al., 2017; Rodler et al., 2018; Wu et al., 2022) modify EVM to obtain more detailed transaction execution information for detecting adversarial transactions.
+BlockGPT (Gai et al., 2023) applies large language model to detect abnormal transactions.
+Qin et al.
+(Qin et al., 2023b) introduce the Execution Property Graph to enable timely detection.
+However, these methods become ineffective in identifying attacks when attackers leverage private mempool services to initiate adversarial transactions.
+Adversarial Contract Detection.
+The study of adversarial contract detection in DeFi attacks remains relatively underexplored.
+Forta (For, 2024) adopts an NLP-based approach for adversarial contract detection; however, its reliance on simple bytecode-level analysis results in limited effectiveness.
+Beyond Forta, the only comparable work is BlockWatchDog (Yang et al., 2024).
+However, their approach is restricted to reentrancy attacks and relies on prior knowledge of victim addresses, thereby limiting its practical utility in real-world scenarios.
+Our method integrates multi-faceted features with semantic tokenization to allow more effective detection across various DeFi attack types.
 9. Conclusion
-In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions.
-We build the first dataset for training PSCFT-based transformers, classifiers and meta classifiers for the binary classification of adversarial contracts.
-Based on this dataset, we found that the use of private mempool services by DeFi attackers is becoming increasingly common, rendering traditional transaction-based detection methods ineffective in most cases.
+In this paper, we propose a new framework for effectively detecting DeFi attacks.
+We build the first large-scale comprehensive dataset for training classifiers for the binary classification of adversarial contracts.
 Through the analysis of DeFi attack patterns, we extract a multi-dimensional feature set along with a special tokenization technique that is able to capture the intrinsic behaviours hidden in adversarial contracts.
-Utilizing these features, we build a transformer, multiple candidate classifiers and a meta classifier and evaluate them comprehensively.
-Experiments show that our method performs exceptionally well in detecting and preventing DeFi attacks with our models reaching F1-Scores as high as $0.8966$ and a false positive rate as low as $0.16\%$.
+Utilizing these features, we carefully design multiple classifiers and evaluate them comprehensively.
+Experiments show that our method performs exceptionally well in detecting adversarial contracts compared to the previous state-of-the-arts with our models reaching F1-Scores as high as $0.8966$ and a false positive rate as low as $0.16\%$ on our labeled datasets.
+10. Data Availability
+To promote open science, we have made our dataset and source code publicly available (loo, 2024).
+Acknowledgements.
+This work is sponsored in part by National Key Research and Development Program of China (2023YFB2704000).